### v1

# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

Abstract
Multimodal large language models (MLLMs) have achieved remarkable performance on complex tasks with multimodal context. However, it is still understudied whether they exhibit modality preference when processing multimodal contexts. To study this question, we first build a MC2 benchmark under controlled evidence conflict scenarios to systematically evaluate modality preference, which is the tendency to favor one modality over another when making decisions based on multimodal conflicting evidence. Our extensive evaluation reveals that all 18 tested MLLMs generally demonstrate clear modality bias, and modality preference can be influenced by external interventions. An in-depth analysis reveals that the preference direction can be captured within the latent representations of MLLMs. Built on this, we propose a probing and steering method based on representation engineering to explicitly control modality preference without additional fine-tuning or carefully crafted prompts. Our method effectively amplifies modality preference toward a desired direction and applies to downstream tasks such as hallucination mitigation and multimodal machine translation, yielding promising improvements. footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
1 Introduction
Multimodal Large Language Models (MLLMs) [3, 41, 43, 50] have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video). Recent advances demonstrate their exceptional capabilities, achieving remarkable performance on complex tasks with multimodal contexts, including autonomous web browsing, graphical user interface understanding, and multimodal dialogue system [16, 18, 40]. Despite impressive performance, fundamental questions remain about their modality preference, particularly whether they solve problems by genuinely integrating cross-modal information or exhibiting modality bias by over-relying on a single dominant modality.
To investigate the modality preferences of MLLMs, a line of research evaluates MLLMs by comparing their performance on unimodal inputs (text-only vs. image-only presentations of the same problem) [12]. However, this approach introduces inherent bias, as it fails to assess the model’s ability to process and integrate truly multimodal inputs where both images and text are simultaneously present. Another line of research analyzes the relative contributions of textual instructions and visual context using Shapley value metric [4, 37, 36]. However, we contend that both parts are indispensable for task resolution, making this framework less adequate for determining true modality preferences across multi-modal contexts.
In this paper, we provide a controlled setup to study the modality preference in MLLMs. As shown in the left panel of Figure 1, we introduce a modality context conflict evaluation setting, where an MLLM is asked to answer a question based on a pair of contrasting evidence. In this way, we can determine the modality preference based on the answer given by MLLMs. Moreover, to dismiss the effect of several confounding factors, including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we utilize widely adopted images and select samples that demonstrate accurate question comprehension and reliable single-modality perception. Building upon this, we introduce a semi-automated annotation framework to construct the refined Modality Context Conflict dataset, MC2, which covers eight recognition-level tasks with 2,000 carefully selected samples. Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 18 representative MLLMs. Our study reveals several intriguing findings:
  • All tested MLLMs generally demonstrate clear modality bias, as illustrated in the right penal of Figure 1, and the modality bias is affected by model-intrinsic factors and task-specific characteristics.
  • The direction of modality preference in MLLMs can be modulated through external interventions, such as modality degradation or instructional guidance.
  • Latent space analysis confirms that modality preferences manifest as geometrically separable patterns in MLLMs’ representations, which can be captured by distinct vectors.
Built on these, we propose a modality preference probing and steering method based on representation engineering to explicitly amplify the direction of modality preference without the need for additional fine-tuning or prompt design. Experimental results show that the proposed method leads to notable performance improvements on both hallucination mitigation and multimodal machine translation. Our main contributions are summarized as follows:
  • We introduce a Modality Context Conflict benchmark for the systematic evaluation of modality preference in MLLMs.
  • Our systematic analysis reveals intrinsic modality preferences in MLLMs, and we find that such preferences are steerable and can be identified through latent representations, offering insights into reasoning under multimodal conflicts.
  • Inspired by these findings, we propose a method to steer modality preference via representation-level intervention, enabling controllable preference adjustment on MC2 and improving performance on downstream tasks.
Figure 1: Illustrations of evaluating modality preference. Left: Use multimodal conflict pairs to evaluate modality preference. Right: Quantified results for modality preferences. Vision and Text represent vision and text preference, other represents irrelevant prediction.
2 Related Work
Multimodal Large language Models.
In recent years, Large Language Models (LLMs) have achieved impressive results across a wide spectrum of tasks [33, 53, 15, 47, 19]. Building upon these, Multimodal Large Language Models (MLLMs) have extended such capabilities, demonstrating strong performance in visual question answering, structured information understanding, and visual navigation [55, 51, 32, 44, 52]. Recent research has turned to systematically evaluating their biases and failure modes [30, 1, 5, 4, 9], aiming to better understand their limitations and guide the development of more reliable models.
Modality Preference in MLLMs. Modality preference—whether MLLMs prioritize one modality over another and to what extent they favor a specific modality when resolving multimodal inputs—can reveal how MLLMs jointly reason over different modalities. Some studies have also observed that MLLMs often favor textual inputs while neglecting visual information, which can result in hallucinations [8, 34, 54]. This highlights the importance of investigating modality preference.
To study the modality preferences of MLLMs, one common approach involves evaluating model performance on unimodal inputs—comparing results when only text or only image inputs are provided for the same task [12]. However, this method inherently overlooks how models handle genuinely multimodal scenarios where both modalities are presented simultaneously, thus limiting its effectiveness in capturing true modality preference. Another research direction assesses the relative influence of textual and visual inputs using Shapley value-based attribution methods [4, 36, 37]. Yet, we argue that both modalities are typically essential for solving multimodal tasks, making this approach insufficient for identifying a model’s inherent modality preference in realistic multimodal contexts. Concurrent work evaluates the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts [45]. However, conflict detection is only one aspect of multimodal reasoning and does not fully reflect a model’s overall ability to jointly reason across modalities.
In this work, in order to investigate their modality preferences, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts. Compared to prior work, we carefully control for potential confounding factors that may influence modality preference, including question understanding, unimodal content recognition, and interference from the model’s internal knowledge. We systematically evaluate modality preference and design a method to controllably steer it. Finally, we demonstrate the effectiveness of our approach on downstream tasks.
Representation Engineering and Activation Editing. Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in activation space [31, 35, 39, 42]. Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs [14, 38, 46, 48, 56]. In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs). Instead of focusing on abstract properties, our method identifies and manipulates activation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
3 The MC2 Benchmark
In this section, we introduce the design and methodology behind the construction of the Multimodal Context Conflict dataset, MC2, intended for evaluating modality preference. We outline the data design philosophy in Section 3.1, followed by the data construction pipeline in Section 3.2 and the evaluation method in Section 3.3.
3.1 Data Design Philosophy
To evaluate modality context preference, we construct modality context conflict pairs that simulate how contexts from different modalities influence model decisions, thereby revealing modality preference. However, the evaluation process may be influenced by various factors such as the internal knowledge within MLLMs, question comprehension and the ability to understand single-modality context. Therefore, on the one hand, we focus on evaluating recognition-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs. On the other hand, we ensure that the evaluated models can accurately answer questions based solely on the context of a single modality, minimizing the influence of confounding factors.
Definitions and Notations: The following definitions and notations are used throughout this paper. In this paper, the image is selected as the vision context,$c^{v}$. We utilize advanced LLMs to generate textual contexts $c^{t}$ that conflict with the vision context $c^{v}$ in relation to the question $q$. $A^{v}$ and $A^{t}$ are answers based on the vision context or textual context.
3.2 Semi-automated Data Construction Pipeline
In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design principles outlined in Section 3.1. The dataset is derived from the TDIUC [22] dataset, sourced from MS-COCO [26], widely adopted in model development to ensure the evaluated models can recognize the images. We extract image caption $caption$, vision context $c^{v}$, question $q$, and answer $A^{v}$ for each sample from TDIUC as the foundation for data annotation. The pipeline follows these steps:
Candidate Textual Context Generation. We use DeepSeekV3 [28] and ChatGPT4o-mini [20] to generate distractor answers $A^{t}$ and candidate textual contexts $c^{t}$ supporting the distractor answers. We aim to construct textual contexts that are aligned with the visual context in terms of overall scene semantics, while introducing content-level conflicts with the visual context specifically in relation to the given question. Therefore, the following task template is employed to generate the distractor answers and textual contexts:
USER: <caption> <question> <answer based on vision context>. Generate the distractor answer different from the given answer based on the vision context and generate a textual context to support the distractor answer. Ensure that the scene described in the textual context is consistent with the caption.
ASSISTANT: <output>
For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection and construction. Detailed prompt templates for different LLMs are provided in Appendix B.1.
Context-based Sample Selecting. We use several basic judge MLLMs, such as LLaVA1.5-7B [29] and QwenVL-7B [6], to select samples. This ensures that all subsequent MLLMs to be evaluated demonstrate high recognition capabilities for both vision and textual contexts. The following task template is used to prompt the judge MLLMs for selecting samples:
USER: <vision/textual context> <question>. Answer question based on the context.
ASSISTANT: <output>
We select only those samples for which all judge MLLMs can correctly answer the given question based on either the vision context or the textual context
Human Verification. Although the previous steps generally yield reliable results, we incorporate manual inspection to ensure the high quality of the data annotation. In this step, we verify the existence of conflicts between $c^{v}$ and $c^{t}$ and ensure that both contexts can correctly direct $q$ to the corresponding answers, $A^{v}$ and $A^{t}$. When errors are found in a sample, we either correct or discard the sample entirely. Each category is cross-verified by three annotators to ensure the reliability of the results. The detailed manual annotation tasks and implementation procedures are provided in Appendix  B.1.
Iterative Refinement. The dataset undergoes multiple rounds of refinement through a feedback loop. This process ensures that the conflicts between different modality contexts are preserved and consistently represented, and MLLMs can correctly understand both the vision and textual contexts. The iterative process helps identify and rectify potential errors, thereby enhancing the dataset quality.
To this end, the dataset is organized into eight categories: positional reasoning, counting and color, attribute recognition, sentiment analysis, activity recognition, sports recognition, and object recognition, with 250 samples in each category. Details of the data annotation format along with sample cases and dataset statistics is presented in Appendix B.3.
3.3 Question Design and Preference evaluation
Question design.
We reformulate the original questions using GPT-4o-mini [21] into binary-choice format. Additionally, to mitigate potential position bias in model outputs, we implement a consistency evaluation strategy. Specifically, for each binary-choice question, we create two versions by swapping the order of the answer choices such as, “How many people are there? A. four. B. five.” and “How many people are there? A. five. B. four.” A model’s response is considered consistent only if it selects different answers in the two versions (i.e., A then B, or B then A). If the model selects the same option in both versions, the response is marked as invalid, indicating possible position bias.
Evaluation for modality preference.
To quantify modality preference, we evaluate whether the model’s prediction is based on visual or textual context according to the following three types:
  • Vision: The prediction aligns with visual context, indicating a preference for vision modality.
  • Text: The prediction aligns with textual context, indicating a preference for text modality.
  • Others: The model gives inconsistent responses or outputs irrelevant answers. These instances are discarded from further analysis.
We use the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/{S_{\text{vision}}+S_{\text{text}}}$. Vision Ratio greater than 0.5 indicates that the model tends to favor visual information over text.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
LLaMAVision 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
LLaVA1.5-13B 34.4/59.6 8.8/89.6 4.8/88.0 12.0/84.8 1.6/96.4 12.8/82.0 9.6/87.6 31.2/62.8 14.4/81.3
OneVision-7B 32.0/36.4 21.6/54.8 2.8/94.4 24.8/56.4 2.4/86.4 30.0/38.0 11.6/71.2 42.4/31.2 20.9/58.6
Owl3-7B 60.8/31.6 16.4/72.4 10.8/85.6 22.0/69.6 8.4/88.0 28.4/60.4 17.2/71.2 60.0/29.6 28.0/63.5
Qwen2VL-7B 26.4/58.0 12.4/82.8 0.8/95.6 13.2/80.4 4.0/93.6 16.0/78.8 11.6/83.6 38.0/54.0 15.3/78.3
Qwen2.5VL-7B 65.6/12.8 45.2/46.0 18.0/68.8 46.4/38.0 51.6/39.6 70.8/20.0 42.0/43.6 77.6/14.0 52.2/35.4
GLM-4V-9B 42.0/42.4 32.4/59.2 8.8/81.6 28.0/62.4 15.2/74.4 56.8/32.8 23.3/66.0 54.0/32.8 32.6/56.5
SPHINX-V2-1K 39.6/50.8 14.8/82.4 1.2/98.4 16.8/77.6 9.2/85.6 23.2/69.2 24.4/67.2 59.2/32.4 23.6/70.5
InternVL3-9B 45.2/35.2 21.2/68.0 20.8/62.4 27.2/54.4 23.2/50.4 38.0/40.4 19.6/63.2 76.8/14.8 34.0/48.6
InternVL3-14B 72.8/8.8 30.8/48.4 25.2/60.0 33.2/52.0 37.2/47.2 58.0/21.2 24.8/52.8 84.4/9.6 45.8/37.5
CogVLM2-19B 44.0/39.6 29.2/56.0 8.8/75.6 19.2/54.8 8.0/73.2 31.6/43.2 25.2/60.8 59.2/28.4 28.2/54.0
LLaVA-next-7B 10.8/74.4 5.2/85.2 0.8/93.2 3.6/79.6 0.4/90.8 6.0/76.0 4.8/73.6 26.0/46.8 7.2/77.5
GPT-4o-mini 94.4/3.2 35.6/47.6 60.4/28.4 22.0/58.9 19.4/59.2 34.8/36.4 71.2/20.4 78.4/12.8 52.0/33.4
Table 1: Results for modality preference. Values are reported as vision/text results for each model.
4 What Modality Preference Has MLLMs Developed?
We use the proposed MC2 dataset to evaluate a wide range of MLLMs, including the proprietary GPT-4o-mini [21] and seventeen open-source MLLMs with varying parameter sizes. For detailed model evaluation, please refer to Appendix C.
We investigate the modality preferences of MLLMs by addressing two key questions: Which modality do MLLMs prioritize? and Can modality preference be controlled? We first analyze the inherent modality preferences of different MLLMs, and then delve into how such preferences can be controlled. This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
4.1 Which Modality Do MLLMs Prioritize?
Before evaluating modality preference, we first assess the ability of MLLMs to answer questions accurately given a single-modality context in the MC2 dataset. All models achieve over 95% accuracy when provided with either textual or visual context. Please refer to Appendix C.1 for detailed results.
Figure 2: Results of modality preference. Left: averaged modality preferences of different MLLMs, Right: vision Ratios for different MLLMs across various tasks.
How do different MLLMs exhibit preferences across various tasks? As described in Section 3.3, we quantify modality preference using model predictions categorized as Vision, Text, and Others, and use the Vision Ratio to quantify the preference for the vision modality, with the results presented in Figure 2 and Table 1. We observe that all MLLMs demonstrate some level of modality preference bias. We observe that the majority of MLLMs exhibit a strong preference for the text modality, particularly as evidenced by the Vision Ratio of LLaVA1.5-13B, which is only 14.4% in the left panel of Figure 2. Interestingly, the Qwen2.5VL and InternVL3 models show a certain degree of preference for the vision modality, which contrasts with the previous conclusion that MLLMs suffer from a severe language prior [23, 37, 45].
A more detailed analysis of different tasks reveals that Qwen2.5VL and InternVL3 have higher Vision Ratios across all tasks compared to other models, especially in high-level recognition tasks such as sports and object recognition, as shown in the right panel of Figure 2. Moreover, we find that low-level tasks (e.g., attribute and color recognition) as well as tasks involving more complex reasoning (such as sentiment analysis, positional reasoning, and counting) consistently exhibit a stronger reliance on the text modality across all MLLMs.
What is the relationship between modality preference and model size? We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference. As shown in the left panel of Figure 3, we observe that for all model families, the preference for the vision modality increases with the model size. And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases. However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their size increases.
Figure 3: Evaluation results of modality preference and attention analyse. The left shows modality preference of MLLMs as parametersize changes, measured by Vision Ratio. A higher Vision Ratio indicates a stronger vision modality preference. the right shows the trends of Vision Ratio and multimodal Attention Ratio across different tasks and models.
4.2 Can Modality Preference Be Controlled?
We employ two intervention methods, content quality manipulation and instruction-based guidance, to explore whether modality preference can be controlled.
Figure 4: Effects of different interventions on the modality preference, measured by Vision Ratio. The left two panels: modifying input quality, including image noise and text grammar errors. The right two panels: using specific instructions to guide MLLMs toward the vision or text modality.
Further analysis using attention attribution. To further investigate the internal mechanisms underlying modality preference, we analyze the attention distributions over the vision and textual contexts in Qwen2.5VL-7B and LLaVA-OneVision-7B. Specifically, we compute the average attention scores for each modality at every token position across all transformer decoder layers of the MLLMs, and define the ratio between visual and textual attention as the Attention Ratio. We then visualize the Attention Ratio alongside the Vision Ratio across a set of representative tasks. As shown in right panel of Figure 3, the trends of Vision Ratio and Attention Ratio are highly aligned across both models and tasks. This suggests that modality preferences are closely related to the internal attention allocation. A similar observation has also been reported in prior work [45].
How does modality content quality affect modality preference? We manipulate the quality of input modality content by introducing Gaussian noise to visual contexts and adding grammatical or spelling errors to textual contexts to observe shifts in modality preference. As shown in the left two panels of Figure 4, we observe that adding noise to images causes a shift in modality preference toward the text modality, whereas introducing grammatical or spelling errors into the text leads to a shift toward the vision modality. It is noteworthy that, despite the degradation in the quality of one modality, the models’ ability to comprehend information from a single modality remains largely unaffected.
This implies that MLLMs may implicitly associate modality reliability with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning. Details of the data generation, experimental design, and full results are provided in Appendix C.2.
Figure 5: PCA visualization of modality preference directions under different intervention strategies. Left: representation shifts under noise-based interventions (image noise or textual errors). Middle: representation shifts under instruction-guided interventions. Right: layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction.
Can modality preference be guided through instruction design? We further investigate the impact of instruction design on modality preference. We design the following task instruction:
USER: <vision context> <textual context> <question>. In this case, there is a conflict between the vision and textual context, and you should rely on the textual/(vision) context rather than the vision/(textual) context.  ASSISTANT: <output>
As shown in the right two panels of Figure 4, instructions directing the model toward either the textual or vision modality effectively guide the modality preference in the specified direction. Moreover, guidance toward the textual modality tends to produce a more pronounced effect compared to guidance toward the vision modality. This tendency may reflect the disproportionate exposure to textual instructions during training, leading to stronger alignment with text-guided cues.
Modality Preference Direction in Representation Space To further understand how the two intervention methods influence modality preference internally, we analyze the hidden representations of the models. Specifically, we apply Principal Component Analysis (PCA) to the hidden states across different layers to identify the dominant direction corresponding to modality preference shifts. Detailed implementation is described in Appendix C.3.
As shown in the left panel of Figure 5, introducing Gaussian noise to visual inputs leads to a noticeable shift in the representation space along a clear direction, while adding grammatical or spelling errors to text inputs results in a comparatively weaker shift. This asymmetry is consistent with our earlier observations in Section 4.2, where textual perturbations steer less pronounced changes in modality preference compared to visual perturbations. Furthermore, the middle panel of Figure 5 illustrates that instruction-based interventions also steer significant representation shifts, aligning with the modality specified in the instruction. Notably, the directionality in the PCA space reveals that the model’s internal states are indeed sensitive to modality control cues, whether they originate from instruction design. Inspired by this, in the following section 5, we propose a probing and Steering method to amplify modality preference and apply it to downstream tasks.
Figure 6: Overall framework of the proposed method. Modality Preference Probing collects the neural activity representations in response to each modality preference prompt. It then computes and scales the direction of modality preference that expresses a preference for a specific modality. Modality Preference Steering selects the target layer during the second inference and adds the scaled modality preference direction to the representation at the corresponding layer at each inference step.
5 Representation-Based Modality Steering
In this section, we adapt the concept of representation engineering [56] to probe the modality preference direction and reveal the consistent representation behavior behind both interventions discussed in Section 4.2. Inspired by this, we steer and amplify the preference direction to control the model’s behavioral expression. Representation engineering involves collecting neural activity by constructing vectors of activation values that cause desired changes to output text, and controlling the model’s behavior by adding the probing vectors to the forward passes of a frozen LLM. As shown in Figure 6, the proposed framework consists of two parts: Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
5.1 Modality Preference Probing
Modality Preference Probing (MPP) is proposed to probe and collect neural activity that represents the direction of modality preference. Inspired by the pre-training objective of decoder-only MLLMs, Next Token Prediction objective [21] and previous works that extract classification features from the last token [11], we collect neural activity from the last token in the input text.
The process involves probing modality preference through two pairs of modality preference requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context rather than the text context’) and another with a text preference probing (e.g., ‘answer the question based on the text context rather than the vision context’). Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$. Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states on the two queries at the last token of the input at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM. We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs. More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards expressing the text modality at layer $l$ for a given query as:
$\mathbf{u}^{t}_{\ell}=\frac{1}{N}\sum_{i}^{N}\big{(}\mathbf{x}_{i,\ell}^{t}-\mathbf{x}_{i,\ell}^{v}\big{)}.$ (1)
Averaging over different queries allows us to capture the activation values most closely associated with modality preference requests, independent of questions.
As shown in the right panel of Figure 5, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples. We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers. Similar patterns are observed for Qwen2VL and LLaVA-OneVision, as detailed in Appendix D.1. Based on this observation, we select the corresponding layers of the model to control the direction of modality preference in Section 5.2.
Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
Vision MLLM-only 26.4 12.4 0.8 13.2 4.0 16.0 11.6 38.0 15.3
Inst Design 60.8 24.0 20.0 20.4 10.8 32.0 27.2 63.2 32.3
CoT Prompting 57.6 27.6 16.0 18.0 21.6 35.6 44.4 52.4 34.2
Few shot 32.0 12.0 10.0 11.0 2.0 25.0 9.0 38.0 17.2
ours 78.8 35.6 38.8 29.6 22.4 56.4 45.2 78.4 48.1
Text MLLM-only 12.8 46.0 68.8 38.0 39.6 20.0 43.6 14.0 35.4
Inst Design 14.4 46.8 72.8 40.4 55.6 24.4 35.6 11.6 37.7
CoT Prompting 27.2 63.6 83.6 62.8 75.6 53.2 58.0 20.4 55.6
Few shot 21.0 77.0 89.0 73.0 73.0 60.0 70.0 42.0 63.1
ours 69.6 67.6 84.4 50.8 82.8 57.6 54.8 41.2 63.6
Table 2: Performance by steering Qwen2VL-7B towards vision and Qwen2.5VL-7B towards text.
5.2 Modality Preference Steering
After probing the direction of the modality preference, we compute the steering vector by re-scaling the direction vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$. The scaling process must carefully balance two objectives: it must be strong enough to effectively steer the model’s modality preference, while simultaneously preserving the model’s normal output behavior. In our preliminary experiments, we observe that setting the weight too large led to repetitive and meaningless outputs, whereas a weight that is too small fails to steer any noticeable change in the model’s modality preference. Unlike previous approaches [56, 38] that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original hidden state distribution. This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities. Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
$\mathbf{s}^{t}_{\ell}=w\mathbf{u}^{t}_{\ell},\ \text{where}\ w=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\mathbf{x}^{t}_{i,\ell}\|}{\|\mathbf{u}^{t}_{\ell}\|}$ (2)
Finally, during a new round of inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ (as identified in Section 5.1) at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens—including vision tokens, text tokens, and newly generated tokens—to steer the model’s response toward the text modality.
Similarly, to steer responses towards the vision modality, the adjustment is performed in the opposite direction. In the final implementation of our method, no additional data or labels are introduced during either the probing or steering stages. We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the model’s modality preference. Notably, the use of unlabeled or self-generated signals avoids annotation bias, which is a critical property when aiming to enhance the generalization ability of different models across diverse domains.
5.3 Verifying the Effectiveness of Steering Modality Preference
We verify the effectiveness of the proposed method in controlling modality preference on MC2 across several MLLMs, including Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-14B. We consider several training-free approaches as baselines: MLLM-only refers to MLLM directly reasoning in modality-conflicting contexts; Inst-Design uses instructions to guide modality preference direction; CoT-prompting enables complex reasoning through intermediate steps; and Few-shot uses a few examples to guide the model in answering. For detailed method implementation and the results of LLaVA-OneVision-7B and InternVL3-14B, refer to Appendix D.2. As shown in Table 2, the proposed method significantly outperforms the baseline approaches, particularly in high-level recognition tasks.
Phd Model Attribute Sentiment Positional Counting Object Avg
Phd-icc MLLM-only 10.0 2.50 3.50 6.0 8.0 6.0
Inst Design 14.5 2.5 1.5 5.5 25.0 9.8
CoT Prompting 5.0 6.0 8.50 6.5 40.5 13.3
Few shot 3.0 0.5 1.5 5.0 2.0 2.4
+ours 10.0 11.0 14.0 5.0 51.5 18.4
Phd-iac MLLM-only 28.5 8.5 20.5 30.5 50.0 27.6
Inst Design 34.5 13.0 26.0 39.0 60.0 34.5
CoT Prompting 15.5 23.5 30.2 17.0 59.0 29.0
Few shot 17.0 9.0 14.5 29.0 37.0 21.3
+ours 29.5 14.5 35.0 28.5 63.5 34.2
Table 3: Performance of the proposed method on Qwen2VL-7B using the hallucination evaluation dataset from Phd [30]. Phd-icc provides an incorrect context in addition to the correct image, while Phd-iac introduces an inaccurate context to assess model robustness under misleading input.
Method en->tr tr->en
Qwen2.5VL-7B 8.92 18.56
      +Inst towards vision 8.21 (-0.71) 16.09 (-2.47)
      +Inst towards Text 9.45 (+0.53) 18.98 (+0.42)
      +ours 10.22 (+1.3) 19.89 (+1.33)
Table 4: The translation results for Ambigcaps using different instruction for preference towards vision modality and text modality and the proposed method steering modality preference towards text modality. We report the Bleu score for the bidirectional translation between English (en) and Turkish (tr).
5.3.1 Downstream Task Applications of Modality Preference Control
We also evaluate the effects of adjusting modality preference towards specific directions on two downstream tasks. Specifically, we assess the hallucination mitigation effect on two subsets of the PhD dataset [30], namely Phd-icc and Phd-iac. Additionally, we evaluate the performance on a multimodal machine translation (MMT) task using the AmbigCaps [25] dataset.
By steering modality preference towards the vision modality, which places greater trust in visual information, we mitigate multimodal hallucinations. As shown in Table 3, our method significantly outperforms the baseline methods on PhD. Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation. This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output. We use our method to steer the modality preference towards the text modality. As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish. For detailed implementation and results, refer to Appendix D.3.
6 Conclusion
This paper investigates modality preference in multimodal large language models (MLLMs). We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference. We observe that all 18 tested MLLMs exhibit overall modality bias, and modality preference directions are encoded in MLLMs’ representation space. Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference. Experiments show that the proposed method generalizes well to downstream tasks, such as hallucination mitigation and multimodal machine translation.
Appendices
All codes, data, and instructions for our MC2 can be found in https://anonymous.4open.science/r/Modality-Preference-8016. MC2 is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
Our supplementary materials are summarized as follows:
Appendix A Discussion
A.1 Limitations
This paper investigates the modality preference in multimodal large language models (MLLMs) using a controlled experiment setup with a modality conflict dataset. In constructing the dataset, we employs LLaVA1.5-7B and QwenVL-7B to filter samples and ensure that most models could answer questions correctly based on a single modality. However, this process requires multiple iterations and turned out to be time-consuming. Therefore, devising a more efficient and elegant method for sample selection may be of greater importance.
A.2 Social Impacts
The proposed MC2 evaluates the modality preference of MLLMs. Understanding which modality a model prioritizes could be used to circumvent safety mechanisms (e.g., hiding harmful content in the favored modality), making it harder for filters to detect inappropriate content. Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
A.3 Ethical Considerations
All images and generated text contexts, which we use to construct MC2, strictly follow guidelines designed to exclude any harmful, unethical, or offensive content. Human annotators are explicitly instructed to refrain from producing any personally identifiable information or inappropriate language during the annotation process. Furthermore, our benchmark does not involve any comparisons of harmful, ethical, or offensive content between image pairs.
A.4 License of Assets
All images in MC2 are publicly available from COCO [26]. We release our benchmark under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
Appendix B Dataset Construction
B.1 Conflict Text Context Generation
Details for data generation using LLMs
To ensure reproducibility and transparency, we include the exact prompts used in our data generation process. These prompts were designed to generate the candidate textual contexts and corresponding answers using GPT-4o-mini [21] and DeepSeekV3 [28]. Below, we provide representative examples of the prompts used during dataset construction given the caption of an image, question, the answer for the question based on image and the task type for the question. For the full list of prompts, please refer to the project repository.
Conflict Context Generation for counting task using DeepSeekV3 Instruction: # Given a description of an image and a corresponding counting type question with its answer, now you are required to generate a text context that points to an answer that fluctuates by 1 or 2 from the original answer. The context explicitly supports the new answer, providing clear evidence that aligns logically with the counting question. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation of for other tasks using DeepSeekV3 Instruction: # Given the caption of an image and a corresponding {task-type} type question with its answer, now you are required to generate a text context as a premise that supports a new distractor answer for the question. The context should mimic the environment described in the caption but should not include {answer based on vision context}, while maintaining logical consistency within the context. Only one alternative answer should be generated. Caption: {caption} Question: {question} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer. The context should support the new answer, include the caption while maintaining logical consistency within the context and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation for other tasks using GPT-4o-mini Instruction: # Given a caption of an image and a corresponding counting question with its answer, you are required to generate a single text context that provides an indirect premise leading to a new answer that fluctuates by 1 or 2 from the original answer. The context should build an indirect premise to the new answer. Carefully design this context. For this task, I want you to first describe the scene with a certain quantity and then introduce an increase or decrease in that quantity to imply the final answer and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Human Verification
Although the text contexts and answers generated by strong LLMs—filtered through judge MLLMs such as LLaVA1.5-7B [29] and QwenVL-7B [6]—generally yield reliable results, we further incorporate manual inspection to ensure the high quality of data annotations. Specifically, we verify that the visual and textual contexts are indeed in conflict, and that each modality independently supports the corresponding answer to the given question. This involves a two-stage manual review process:
  • Modality-Answer Alignment. First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question. This step is particularly important because the original VQA answers in the TDIUC [22] dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
  • Conflict Verification. Next, annotators examine whether the visual and textual contexts are semantically inconsistent with respect to the question. That is, the two modalities should lead to different correct answers when considered separately. Samples where both modalities lead to the same answer are discarded, as they do not reflect a true modality conflict.
Samples that do not meet either verification criterion are flagged for further review. Depending on the nature and severity of the issue, we take one of the following actions: revise the prompt to improve clarity, regenerate the problematic part of the sample (e.g., the question or context), or discard the sample entirely if it cannot be reasonably corrected.
To ensure consistency and reduce subjectivity, each category (i.e., vision-aligned, text-aligned, and conflict) is independently verified by three trained annotators. Disagreements are resolved through discussion or majority voting. In addition, we conduct random spot-checks throughout the dataset to ensure the consistency and reliability of the annotations.
B.2 Data Statistics
We computed the average number of words in the text context for all samples within each task type using the spaCy library. [^0] We use the spaCy library in Python, available at https://pypi.org/project/spacy/. As shown in Table 5, while there are some variations in text length across tasks, the differences are relatively minor. This indicates that text length is unlikely to be a confounding factor in evaluating modality preference across different task types.
B.3 Illustrative Samples from the MC2 Benchmark
To provide an intuitive understanding of the MC2 benchmark and the nature of modality conflict, we present a few representative samples covering different task types as shown in Figure 7, Figure 8, Figure 9 and Figure 10.
<image> is a placeholder for below image
User: <image> Conflict Text Context: Three sheep are peacefully eating grass, surrounded by lush greenery. Their heads are lowered as they nibble on the fresh blades, completely undisturbed. Question: What are the cows in the back doing?
Assistant: <output>
vision-based Answer: running
Text-based Answer: eating
<image> is a placeholder for below image
User: <image> Conflict Text Context: In the photo, there are three boys playing Frisbee, and one more boy is partially visible in the corner, bending down to tie his shoelaces, making a total of four people. Question: How many people are in the photo?
Assistant: <output>
vision-based Answer: five
Text-based Answer: four
Figure 7: Illustration of using modality context conflict pairs to investigate modality preference in activity recognition (Left) and counting tasks (Right). The highlighted areas indicate the points of conflict between visual and textual contexts.
<image> is a placeholder for below image
User: <image> Conflict Text Context: The birthday cake was designed to look like a sleek police car, complete with edible flashing lights and a fondant badge on the side. Question: What is the cake in the shape of?
Assistant: <output>
vision-based Answer: fire truck
Text-based Answer: police car
<image> is a placeholder for below image
User: <image> Conflict Text Context: Two wildebeests are standing in a dry, grass-less savanna, their dark coats contrasting with the dusty ground. The area is sparse, with only a few scattered shrubs visible in the background. Question: What animal is shown?
Assistant: <output>
vision-based Answer: zebras
Text-based Answer: wildebeests
Figure 8: Illustration of using modality context conflict pairs to investigate modality preference in attribute recognition (Left) and object recognition tasks (Right). The highlighted areas indicate the points of conflict between visual and textual contexts.
<image> is a placeholder for below image
User: <image> Conflict Text Context: A large brown clock tower mounted in the face ofa building overlooks a vibrant park filled with lush green trees. The contrast between the brown tower and the surrounding greenery creates a picturesque scene. Question: What color are the trees?
Assistant: <output>
vision-based Answer: white
Text-based Answer: green
<image> is a placeholder for below image
User: <image> Conflict Text Context: A white bus with a large rack on the front is parked by the beach, designed to carry equipment for surfing .The rack is sturdy and spacious, perfect for securing bulky items. Question: What can you hang on the rack on the front of the bus?
Assistant: <output>
vision-based Answer: bikes
Text-based Answer: surfboards
Figure 9: Illustration of using modality context conflict pairs to investigate modality preference in color recognition (Left) and positional reasoning (Right) tasks. The highlighted areas indicate the points of conflict between visual and textual contexts.
<image> is a placeholder for below image
User: <image> Conflict Text Context: A girl sitting at a counter with a piece of pizza, staring blankly at the wall while the pizza grows cold in front of her. The room is quiet, and she seems uninterested in her surroundings. Question: What is the girl on the right feeling in the image?
Assistant: <output>
vision-based Answer: happy
Text-based Answer: bored
<image> is a placeholder for below image
User: <image> Conflict Text Context: The young girl is running swiftly across the field, dribbling a soccer ball with precision as she maneuvers past imaginary opponents. Her focus is on scoring a goal, and she practices her footwork with determination. Question: What sport is depicted in the picture?
Assistant: <output>
vision-based Answer: tennis
Text-based Answer: soccer
Figure 10: Illustration of using modality context conflict pairs to investigate modality preference in sentiment understanding and object recognition tasks. The highlighted areas indicate the points of conflict between visual and textual contexts.
statistics Sport Attribute Sentiment Positional Counting Color Activity Object Avg
Text Length 52.48 33.50 39.69 31.53 37.12 31.15 49.68 39.71 39.36
Table 5: Average text context length across different task types in the MC2 dataset.
Appendix C Model Evaluatiion
C.1 Evaluatiion Detail for Modality Preference
We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B [29], LLama3.2-11B-Vision-Instruct [13], LLaVA-OneVision-7B/72B [24], CogVLM2-19B [17], mPLUG-Owl3-24-07 [49], Qwen2VL-7B [43], GLM-4V-9B [10], SPHINX-V2-1K [27], InternVL3-9B/14B/38B/78B [2], LLaVA-next-7B/13B/34B [29] and Qwen2.5VL-7B/32B/72B [7]. All the open-source models are evaluated using NVIDIA A100 or A800 GPUs. We also evaluate the proprietary model, GPT-4o-mini [21] via the official API.
Details of single-modality context evaluation
Before evaluating modality preference, we first assess the ability of MLLMs to answer questions accurately given a single-modality context in the MC2 dataset. Specifically, we evaluate the models’ accuracy in answering based on text context and based on vision context (based on the image). As shown in Table 10 and Table 11, all models achieve over 95% accuracy when provided with either textual or visual context. This indicates that question understanding and the understanding of single-modality context do not affect the modality preference evaluation. Therefore, we have excluded this confounding factor from the analysis.
Details of results for modality preference evaluation
We provide the results of modality preference for several models in Table 1 in the main text. More detailed modality preference evaluation results are presented in Table 7. We evaluated 18 models, including 17 open-source models and the commercial model, GPT4o-mini.
C.2 The Details for Controlling Modality Preference
Noisy image and text grammar error generation
For noisy images, we initially adopte two approaches: adding Gaussian noise to the image and applying a uniform linear motion blur kernel. In our experiments, both methods yield similar results in terms of model performance, indicating comparable effects on the evaluation. We report the results based on the Gaussian noise method in this paper.
To generate grammatically incorrect text, we employ DeepSeek-V3 to rewrite the original text context. We find that DeepSeekV3 performs well in introducing localized grammatical errors while preserving the overall structure. We use two different prompting strategies, both of which produced similar results. In this paper, we report the results based on the V1 version of grammar error generation. The specific prompt used is shown below:
Text Context Grammar Error Generation (V1) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes. These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words. Despite these errors, the core meaning of the sentence must remain unchanged. The text to be rewritten is: {original text context}. Output the revised text enclosed in <context2> </context2> tag. Only output the revised text and don’t explain anything.
Text Context Grammar Error Generation (V2) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes. These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words. Despite these errors, the core meaning of the sentence must remain unchanged. Additionally, do not alter the meaning related to the provided critical word. The text to be rewritten is: {original text context}. the critical word is: {answer based on text context}. Output the revised text enclosed in <context1> </context1> tag. Only output the revised text and don’t explain anything.
The detail for controlling modality preference using noise data
We first test the model’s ability to understand a single noisy modality (either a noisy image context or a text context with grammar errors). As shown in Table 13 and Table 12, the model’s understanding of the single-modality content remains close to 95% for the noisy image context and close to 98% for the noisy text context, which is nearly the same as its performance on the original single-modality content (shown in Table 11 and Table 10). This rules out the possibility that the shift in modality preference due to the added noise is caused by a decrease in the model’s understanding of the single modality. Instead, it suggests that MLLMs may implicitly associate modality preference with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning.
We provide the results of modality preference using either the noisy image context or the text context with grammer error in Figure 4 in main text. We also provide more results for more models in Table 14 and Table 15.
More results for controlling modality preference through instruction design.
In Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B and Qwen2.5VL-7B. We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 8 and Table 9. For each setting, we report the probability of responses exhibiting a text preference (measured by vision-based accuracy) and the probability of responses exhibiting a vision preference (measured by text-based accuracy).
C.3 The details of PCA Analysis
In Section 4.2, we mentioned the PCA analysis details regarding the "Modality Preference Direction in Representation Space." Here, we provide a more in-depth description of the setup. We extract the model’s hidden representations from the last token of the input across different layers. Then, we apply the PCA method to reduce the dimensionality to two dimensions for visualization. The following settings were visualized:
  1. 1.
    The model states under the original modality context input in conflicting scenarios.
  2. 2.
    The model states when there is image noise or textual syntax errors.
  3. 3.
    The model states when specific instructions biased towards image or text are added.
To improve PCA dimensionality reduction efficiency, we selected 500 samples for each setting. Additionally, we calculated the center position after dimensionality reduction for each setting. The center (or centroid) of the samples is computed by taking the mean of the reduced-dimensional points across all the samples.
Appendix D Method Applying
D.1 Details for Pattern of hidden states
In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B. As shown in Figure 11, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-14B.
For each model, we selected layers with large absolute differences and small standard deviations. This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
Figure 11: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B models from left to right.
D.2 More Results for Steering Models Toward Vision or Text Modality Preference
In the main text, we present the results of steering Qwen2VL-7B and Qwen2.5VL-7B towards vision and text modality preferences, respectively. Here, we provide additional results on modality preference control, including experiments with InternVL3-14B and LLaVA-OneVision-7B. As shown in Table 6, the proposed method significantly outperforms the baseline in both directions of modality preference.
D.3 Downstream Task Applications
By steering the model’s modality preference towards the vision modality, we reduce vision hallucinations in MLLMs. Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation. This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output.
Details of Vision Hallucination Mitigation
PhD [30] is a multimodal hallucination evaluation benchmark that increases the risk of hallucination by introducing "hitem" items, intentionally guiding the model to produce hallucinated words, thereby increasing the difficulty of the model’s responses. Phd-icc and Phd-ica are subsets of PhD, which respectively increase hallucination risks by introducing incorrect text context and inaccurate text context. For testing convenience, we randomly selected 1,000 samples from the original Phd-cc and Phd-ica datasets for evaluation.
By steering modality preference towards the vision context, the model is encouraged to focus more on visual information, which helps reduce vision hallucinations. As shown in Table 3, the proposed method outperforms most baseline methods across nearly all subcategories of the Phd-icc and Phd-ica datasets. Furthermore, a detailed analysis of the two datasets reveals that most samples in Phd-iac introduce only irrelevant text context, whereas Phd-icc introduces incorrect text context, which makes it easier for the model to generate hallucinations. Proposed method significantly outperforms the baseline methods on Phd-icc, indicating that the proposed method is more effective in mitigating hallucinations.
Details of Multimodal Multimodal machine translation
Ambigcaps [25] benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets. We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality. We also use our method to steer the modality preference towards the text modality. As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish.
Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
Vision MLLM-only 32.0 21.6 2.8 24.8 2.4 30.0 11.6 42.4 20.9
Inst Design 55.6 31.2 12.0 30.8 3.2 36.4 22.4 61.2 31.6
ours 75.6 40.8 18.0 38.0 10.0 53.2 32.0 74.4 42.8
Text MLLM-only 8.8 48.4 60.0 52.0 47.2 21.2 52.8 9.6 37.5
Inst Design 44.8 68.4 82.4 54.8 43.2 50.0 58.8 19.6 52.8
ours 49.6 71.2 85.2 58.0 47.6 48.4 68.4 58.4 57.1
Table 6: Performance by steering LLaVA-OneVision-7B towards vision and InternVL3-7B towards text.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
LLaMAVision 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
LLaVA1.5-7B 20.0/59.6 8.0/88.0 2.0/86.0 8.8/75.2 1.2/96.0 10.8/82.0 9.6/78.8 35.2/52.0 11.9/77.2
LLaVA1.5-13B 34.4/59.6 8.8/89.6 4.8/88.0 12.0/84.8 1.6/96.4 12.8/82.0 9.6/87.6 31.2/62.8 14.4/81.3
OneVision-7B 32.0/36.4 21.6/54.8 2.8/94.4 24.8/56.4 2.4/86.4 30.0/38.0 11.6/71.2 42.4/31.2 20.9/58.6
Owl3-24-07 60.8/31.6 16.4/72.4 10.8/85.6 22.0/69.6 8.4/88.0 28.4/60.4 17.2/71.2 60.0/29.6 28.0/63.5
QwenVL-7B 35.2/43.6 15.2/63.2 11.2/78.0 16.4/45.6 2.8/73.6 34.0/40.8 19.2/61.6 35.2/31.2 21.1/54.7
Qwen2VL-7B 26.4/58.0 12.4/82.8 0.8/95.6 13.2/80.4 4.0/93.6 16.0/78.8 11.6/83.6 38.0/54.0 15.3/78.3
Qwen2.5VL-7B 65.6/12.8 45.2/46.0 18.0/68.8 46.4/38.0 51.6/39.6 70.8/20.0 42.0/43.6 77.6/14.0 52.2/35.4
InternLM-XC2.5-7B 72.0/19.6 26.4/63.2 50.1/36.4 22.0/65.2 9.2/83.2 50.4/38.0 47.6/41.2 68.8/21.6 43.4/46.1
GLM-4V-9B 42.0/42.4 32.4/59.2 8.8/81.6 28.0/62.4 15.2/74.4 56.8/32.8 23.3/66.0 54.0/32.8 32.6/56.5
SPHINX-V2-1K 39.6/50.8 14.8/82.4 1.2/98.4 16.8/77.6 9.2/85.6 23.2/69.2 24.4/67.2 59.2/32.4 23.6/70.5
InternVL3-9B 45.2/35.2 21.2/68.0 20.8/62.4 27.2/54.4 23.2/50.4 38.0/40.4 19.6/63.2 76.8/14.8 34.0/48.6
InternVL3-14B 72.8/8.8 30.8/48.4 25.2/60.0 33.2/52.0 37.2/47.2 58.0/21.2 24.8/52.8 84.4/9.6 45.8/37.5
CogVLM2-19B 44.0/39.6 29.2/56.0 8.8/75.6 19.2/54.8 8.0/73.2 31.6/43.2 25.2/60.8 59.2/28.4 28.2/54.0
InternVL3-38B 75.2/9.6 45.2/33.6 19.6/60.8 44.0/42.0 41.6/40.0 48.4/29.6 50.4/23.2 84.4/8.0 51.1/30.8
InternVL3-78B 92.4/3.2 46.0/28.8 66.4/18.4 41.6/37.2 69.6/13.2 76.4/8.8 74.4/12.8 89.6/4.0 69.5/15.8
Qwen2.5-VL-32B 85.60/10.40 49.20/39.20 49.60/42.80 52/37.60 52/42 70.80/20 57.20/35.20 86.80/10.40 62.90/29.70
Qwen2.5VL-72B 93.6/4.4 59.2/27.2 50.0/41.2 73.6/19.2 63.6/29.2 83.6/9.6 74.0/21.2 89.2/8.0 73.4/20.0
OneVision-72B 47.2/46.0 20.0/70.8 4.0/93.6 22.8/67.2 12.8/83.2 21.6/60.8 20.8/70.8 71.6/21.2 27.6/64.2
LLaVA1.6-7B 10.8/74.4 5.2/85.2 0.8/93.2 3.6/79.6 0.4/90.8 6.0/76.0 4.8/73.6 26.0/46.8 7.2/77.5
LLaVA1.6-13B 16.0/66.4 7.2/90.4 0.8/92.0 6.4/91.6 2.4/95.6 6.8/88.0 10.0/84.4 22.4/63.2 9.0/84.0
LLaVA1.6-34B 34.8/42.4 12.0/81.6 6.8/85.6 16.8/76.0 11.2/83.2 25.2/60.8 14.0/76.0 60.0/31.6 22.6/67.2
GPT-4o-mini 94.4/3.2 35.6/47.6 60.4/28.4 22.0/58.9 19.4/59.2 34.8/36.4 71.2/20.4 78.4/12.8 52.0/33.4
Table 7: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided but the instruction does not specify which modality context should be used. Values are reported as vision-based accuracy/text-based accuracy for each model.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 55.6/16.4 31.2/37.2 12.0/76.8 30.8/42.4 3.2/77.6 36.4/18.4 22.4/47.6 61.2/16.4 31.6/41.6
Qwen2VL-7B 60.8/26.8 24.0/69.2 20.0/69.6 20.4/74.0 10.8/80.0 32.0/52.0 27.2/61.6 63.2/28.8 32.3/57.8
Qwen2.5VL-7B 77.6/14.4 43.2/46.8 18.4/72.8 43.2/40.4 35.6/55.6 58.8/24.4 53.6/35.6 81.2/11.6 51.4/37.7
CogVLM2-19B 73.2/13.2 47.6/32.4 35.6/28.4 26.8/45.2 14.0/40.0 61.6/17.6 56.0/28.0 76.0/15.2 48.9/27.5
InternLM-XC2.5-7B 84.0/9.6 46.4/42.8 74.0/18.4 36.0/52.4 22.8/66.0 63.6/20.8 74.0/18.4 76.4/15.6 59.7/30.5
GLM-4V-9B 75.2/18.4 48.8/39.6 28.8/54.0 33.6/55.6 38.4/54.0 76.4/16.4 48.4/38.8 80.0/11.6 53.7/36.1
SPHINX-V2-1K 52.4/38.4 16.4/78.8 2.0/97.2 20.8/72.8 13.6/80.8 30.0/58.8 40.8/52.8 64.8/29.2 30.1/63.6
InternVL3-9B 96.0/2.0 67.2/18.8 82.8/13.2 54.8/26.4 55.6/21.2 84.4/7.6 82.8/6.4 91.6/4.0 76.9/12.4
InternVL3-14B 98.4/0.8 86.0/4.4 87.6/7.6 71.6/12.8 78.0/6.8 97.2/0.8 90.8/3.2 96.4/1.6 88.2/4.8
LLaVA1.6-7B 33.2/54.0 6.8/80.8 6.0/82.4 6.4/79.6 2.8/90.4 10.8/70.0 13.6/69.2 48.4/40.8 16.0/70.9
LLaVA1.6-13B 41.6/40.4 10.4/85.2 4.0/62.8 8.4/83.6 5.6/92.8 14.4/70.8 24.8/58.0 45.2/41.2 19.3/66.9
LLaVA1.6-34B 84.8/12.0 48.0/36.4 62.8/24.0 34.0/52.0 38.8/44.4 76.4/14.4 62.0/18.4 80.4/12.4 60.9/26.8
Table 8: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on visual modality context. Values are reported as vision-based accuracy/text-based accuracy for each model.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 45.6/16.4 22.4/37.2 5.6/76.8 27.2/42.4 1.6/77.6 28.8/18.4 16.8/47.6 56.0/16.4 25.5/41.6
Qwen2VL-7B 51.6/34.8 14.8/78.4 6.8/88.0 15.6/79.6 4.0/90.8 18.0/70.8 19.2/72.8 57.6/36.0 23.4/68.9
Qwen2.5VL-7B 77.6/14.4 43.2/46.8 18.4/72.8 43.2/40.4 35.6/55.6 58.8/24.4 53.6/35.6 81.2/11.6 51.4/37.7
CogVLM2-19B 53.6/29.6 28.4/47.6 10.4/56.8 17.6/56.8 6.0/62.0 36.0/35.2 34.0/39.6 67.2/19.6 31.6/43.4
InternLM-XC2.5-7B 79.2/16.0 31.2/56.8 60.4/29.2 26.0/62.8 11.6/75.6 52.8/29.2 60.8/29.6 72.0/18.0 49.2/39.7
GLM-4V-9B 53.2/32.8 30.4/61.2 6.0/85.6 23.2/68.0 20.0/70.0 52.4/35.6 28.0/60.8 68.0/22.0 35.2/54.5
SPHINX-V2-1K 48.4/41.2 14.4/81.6 2.0/98.0 19.2/77.6 11.2/84.0 27.2/67.2 30.8/65.2 63.2/30.8 27.1/68.2
InternVL3-9B 41.2/27.2 13.6/71.6 22.4/60.8 16.8/64.0 18.0/60.4 25.6/46.4 29.2/49.2 62.4/17.6 28.6/49.6
InternVL3-14B 28.4/44.8 14.0/68.4 3.6/82.4 21.2/54.8 28.0/43.2 24.8/50.0 17.2/58.8 55.2/19.6 24.0/52.8
Table 9: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on the textual modality. Values are reported as vision-based accuracy/text-based accuracy for each model.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
LLaMAVision 97.6 97.2 99.6 99.2 97.2 96.0 97.6 97.6 97.8
LLaVA1.5-7B 98.0 98.0 100.0 97.6 98.4 99.2 97.6 97.6 98.3
LLaVA1.5-13B 97.2 97.6 99.6 97.6 97.6 98.8 95.2 99.2 97.9
OneVision-7B 98.0 95.2 100.0 98.4 98.0 98.8 98.0 100.0 98.3
Owl3 97.6 97.2 99.6 98.8 98.8 99.2 99.2 100.0 98.8
Qwen2VL-7B 98.8 96.4 99.6 99.6 98.8 100.0 98.8 100.0 99.0
Qwen2.5VL-7B 99.2 97.6 100.0 99.6 96.8 98.8 98.4 99.2 98.7
CogVLM2-19B 98.0 95.2 99.2 96.0 94.8 98.4 98.0 99.6 97.4
InternLM-XC2.5-7B 96.4 97.6 99.2 99.2 98.0 98.0 98.8 98.4 98.2
GLM-4V-9B 98.4 95.2 99.6 97.2 98.8 98.4 99.6 99.6 98.4
SPHINX-V2-1K 98.4 97.6 99.2 98.8 98.0 99.2 98.4 99.6 98.7
InternVL3-9B 97.6 98.0 99.6 99.2 95.6 96.8 98.8 99.2 98.1
InternVL3-14B 98.4 98.4 100.0 99.2 95.6 98.4 98.8 99.6 98.5
InternVL3-38B 97.6 96.8 100.0 98.8 96.0 97.2 98.4 100.0 98.1
InternVL3-78B 97.2 97.6 100.0 98.0 96.4 96.8 98.0 100.0 98.0
Qwen2.5VL-72B 99.6 98.4 96.8 100.0 97.2 100.0 99.6 99.2 98.9
OneVision-72B 100.0 97.6 97.6 99.6 96.4 100.0 100.0 98.8 98.7
GPT-4o-mini 97.6 97.2 99.6 98.6 97.4 98.4 98.4 100.0 98.4
Table 10: Accuracy of question answering in the MC2 dataset when only unimodal textual context is provided.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
LLaMAVision 100.0 98.8 92.8 98.4 96.4 99.2 98.8 97.2 97.7
LLaVA1.5-7B 99.6 98.0 96.4 100.0 97.6 99.6 98.8 98.4 98.5
LLaVA1.5-13B 99.6 95.2 94.4 97.6 95.2 98.4 96.4 98.4 96.9
OneVision-7B 100.0 97.2 97.2 98.4 84.4 99.6 97.2 98.8 96.6
Owl3 99.2 94.0 94.0 97.2 88.4 96.8 97.2 99.2 95.8
QwenVL-7B 100.0 98.8 97.6 100.0 99.6 100.0 100.0 98.4 99.3
Qwen2VL-7B 99.6 98.8 95.6 98.4 96.4 100.0 99.6 98.4 98.3
Qwen2.5VL-7B 99.6 98.8 98.0 100.0 99.2 100.0 100.0 98.8 99.3
CogVLM2-19B 99.6 99.2 91.2 96.8 91.6 98.8 98.4 98.8 96.8
InternLM-XC2.5-7B 98.8 98.8 94.0 98.0 94.8 99.6 100.0 98.8 97.9
GLM-4V-9B 99.6 99.2 98.0 99.2 97.6 100.0 99.2 99.6 99.1
SPHINX-V2-1K 98.8 97.6 99.2 92.8 98.0 99.6 96.8 99.2 97.8
InternVL3-9B 98.8 95.6 95.6 96.8 90.0 100.0 98.0 98.0 96.6
InternVL3-14B 99.2 96.4 96.4 98.4 92.4 98.8 97.2 98.4 97.1
InternVL3-38B 100.0 98.0 97.2 100.0 94.4 99.6 99.2 98.8 98.4
InternVL3-78B 99.2 99.6 96.8 98.8 96.0 100.0 99.2 98.4 98.5
Qwen2.5VL-72B 97.2 97.2 100.0 99.2 97.2 98.4 98.4 99.6 98.4
OneVision-72B 100.0 97.6 97.6 99.6 96.4 100.0 100.0 98.8 98.7
GPT-4o-mini 100.0 92.0 95.6 100.0 100.0 96.0 96.4 96.0 97.0
Table 11: Accuracy of question answering in the MC2 dataset when only unimodal visual context is provided.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 98.0 95.2 100.0 97.2 97.2 97.2 98.0 99.2 97.8
Qwen2.5VL-7B 98.8 96.4 100.0 99.2 97.6 97.6 98.8 99.2 98.4
CogVLM2-19B 98.0 94.4 99.2 96.0 91.2 97.6 97.6 99.6 96.7
InternLM-XC2.5-7B 97.6 97.2 99.2 99.2 97.6 97.6 98.8 98.8 98.2
GLM-4V-9B 98.8 96.4 99.6 97.2 98.8 98.4 99.6 100.0 98.6
SPHINX-V2-1K 98.4 97.6 99.2 98.8 97.6 98.4 98.4 99.6 98.5
LLaVA1.6-7B 98.0 90.8 98.8 96.8 96.8 97.2 96.0 96.4 96.4
Table 12: Accuracy of question answering in the MC2 dataset when only unimodal noisy textual context is provided.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 98.4 95.6 94.0 98.4 78.4 99.2 95.6 98.0 94.7
Qwen2.5VL-7B 98.8 96.4 100.0 99.2 99.2 100.0 100.0 98.8 99.3
CogVLM2-19B 99.6 97.6 87.6 94.4 91.6 98.8 97.2 98.48 95.7
InternLM-XC2.5-7B 98.4 96.0 94.4 94.0 95.2 100.0 97.6 95.2 96.4
GLM-4V-9B 99.2 98.4 95.6 98.4 98.4 99.6 100.0 99.2 98.6
SPHINX-V2-1K 98.4 97.2 90.0 98.8 98.4 97.6 96.8 98.0 96.9
LLaVA1.6-7B 98.4 94.8 92.4 93.6 84.0 96.8 93.6 95.6 93.7
Table 13: Accuracy of question answering in the MC2 dataset when only unimodal noisy visual context is provided.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 42/29.2 31.2/41.6 4.4/88 31.6/44 4.4/82 39.2/29.6 16.4/66.8 50.8/26 27.5/50.9
Qwen2.5VL-7B 74.0/12.0 53.20/36.40 24.4/62.4 58/29.2 54.8/35.2 78/16 52.8/35.6 82.4/12.4 59.7/29.9
InternLM-XC2.5-7B 75.6/18.4 32.8/57.2 56.0/36.0 26.4/59.2 11.6/78.0 61.2/28.0 52.8/36.4 74.0/20.0 48.8/41.7
GLM-4V-9B 48.0/31.6 37.6/48.8 10.0/81.2 34.0/56.4 18.8/70.8 64.8/27.2 33.6/55.6 61.6/25.2 38.6/49.6
SPHINX-V2-1K 46.0/44.8 18.0/77.6 1.2/98.0 21.6/70.8 11.6/80.8 30.0/64.8 30.8/60.0 62.8/28.8 27.8/65.7
CogVLM2-19B 46.0/39.6 34.0/47.2 12.4/68.4 22.8/52.0 8.4/70.0 37.2/39.2 30.0/55.2 62.8/24.4 31.7/49.5
LLaVA1.6-7B 11.2/71.2 6.4/83.2 1.2/91.6 5.2/76.4 0.4/88.8 7.2/71.6 5.6/70.4 28.8/44.4 8.3/74.7
Table 14: Accuracy of question answering in the MC2 dataset when both noisy textual context and visual context are provided but the instruction does not specify which modality context should be used. Values are reported as vision-based accuracy/text-based accuracy for each model.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
OneVision-7B 24/44 16.4/66.4 2/96.8 17.2/66.4 0.80/87.2 20.4/47.2 8/78.8 35.6/42.8 15.5/66.2
Qwen2.5VL-7B 48.8/23.6 25.6/64.8 6/88 33.2/59.6 39.2/54.8 58.8/32.8 30/56.8 68/56.8 38.7/50.3
InternLM-XC2.5-7B 52.8/33.6 15.2/76.8 24.4/64.4 12.8/78.4 3.6/90.4 33.2/56.8 24.4/64.8 54.8/37.2 27.7/62.8
GLM-4V-9B 32.4/50.4 19.6/73.2 4.0/91.6 21.2/69.2 11.2/78.4 41.6/46.8 15.2/76.8 41.2/42.0 23.3/66.1
SPHINX-V2-1K 29.2/58.0 11.2/86.4 0.8/98.8 16.8/80.4 5.6/90.4 18.0/76.4 18.0/76.0 52.8/40.4 19.1/75.9
CogVLM2-19B 32.0/45.2 22.0/60.0 3.6/84.8 14.0/60.8 6.4/73.2 18.8/49.2 18.8/65.6 46.0/36.4 20.2/59.4
LLaVA1.6-7B 8.4/79.2 4.0/87.2 0.4/94.8 3.2/82.4 0.4/92.0 4.4/77.6 4.4/76.8 22.8/50.8 6.0/80.1
Table 15: Accuracy of question answering in the MC2 dataset when both textual and noisy visual contexts are provided but the instruction does not specify which modality context should be used. Values are reported as vision-based accuracy/text-based accuracy for each model.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. 2.
    Limitations
  5. Question: Does the paper discuss the limitations of the work performed by the authors?
  6. Answer: [Yes] See discussions in Suppl.
  7. 3.
    Theory assumptions and proofs
  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  9. Answer: [N/A]
  10. 4.
    Experimental result reproducibility
  11. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  12. Answer: [Yes] See details in Suppl.
  13. 5.
    Open access to data and code
  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  15. Answer: [Yes] See details in Suppl.
  16. 6.
    Experimental setting/details
  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  18. Answer: [Yes] See details in Suppl.
  19. 7.
    Experiment statistical significance
  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  21. Answer: [Yes] See details in Suppl.
  22. 8.
    Experiments compute resources
  23. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  24. Answer: [Yes] See details in Suppl.
  25. 9.
    Code of ethics
  26. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  27. Answer: [Yes]
  28. 10.
    Broader impacts
  29. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  30. Answer: [Yes]
  31. 11.
    Safeguards
  32. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  33. Answer: [Yes] See details in Suppl.
  34. 12.
    Licenses for existing assets
  35. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  36. Answer: [Yes]
  37. 13.
    New assets
  38. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  39. Answer: [Yes] See detail in Suppl.
  40. 14.
    Crowdsourcing and research with human subjects
  41. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  42. Answer: [Yes] See details in Suppl.
  43. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  44. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  45. Answer: [Yes] See details in Suppl.
  46. 16.
    Declaration of LLM usage
  47. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  48. Answer: [Yes] See details in Suppl.

### v2

@@
 Abstract
-Multimodal large language models (MLLMs) have achieved remarkable performance on complex tasks with multimodal context.
-However, it is still understudied whether they exhibit modality preference when processing multimodal contexts.
-To study this question, we first build a MC2 benchmark under controlled evidence conflict scenarios to systematically evaluate modality preference, which is the tendency to favor one modality over another when making decisions based on multimodal conflicting evidence.
-Our extensive evaluation reveals that all 18 tested MLLMs generally demonstrate clear modality bias, and modality preference can be influenced by external interventions.
-An in-depth analysis reveals that the preference direction can be captured within the latent representations of MLLMs.
-Built on this, we propose a probing and steering method based on representation engineering to explicitly control modality preference without additional fine-tuning or carefully crafted prompts.
-Our method effectively amplifies modality preference toward a desired direction and applies to downstream tasks such as hallucination mitigation and multimodal machine translation, yielding promising improvements.
+Multimodal large language models (MLLMs) have achieved remarkable success on complex multimodal tasks.
+However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multimodal contexts.
+To study this question, we introduce MC2 benchmark, which constructs controlled evidence-conflict scenarios to systematically evaluate modality preference in decision-making.
+Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performances of MLLMs.
+Further analysis shows that modality preference can be controlled by instruction guidance and captured within the latent representations of MLLMs.
+Built on these insights, we propose a probing and steering method based on representation engineering to explicitly control modality preference without requiring additional fine-tuning.
+This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple downstream applications, including multimodal visual understanding and multimodal machine translation.
 footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
 1 Introduction
-Multimodal Large Language Models (MLLMs) [3, 41, 43, 50] have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video).
-Recent advances demonstrate their exceptional capabilities, achieving remarkable performance on complex tasks with multimodal contexts, including autonomous web browsing, graphical user interface understanding, and multimodal dialogue system [16, 18, 40].
-Despite impressive performance, fundamental questions remain about their modality preference, particularly whether they solve problems by genuinely integrating cross-modal information or exhibiting modality bias by over-relying on a single dominant modality.
-To investigate the modality preferences of MLLMs, a line of research evaluates MLLMs by comparing their performance on unimodal inputs (text-only vs.
-image-only presentations of the same problem) [12].
-However, this approach introduces inherent bias, as it fails to assess the model’s ability to process and integrate truly multimodal inputs where both images and text are simultaneously present.
-Another line of research analyzes the relative contributions of textual instructions and visual context using Shapley value metric [4, 37, 36].
-However, we contend that both parts are indispensable for task resolution, making this framework less adequate for determining true modality preferences across multi-modal contexts.
-In this paper, we provide a controlled setup to study the modality preference in MLLMs.
-As shown in the left panel of Figure 1, we introduce a modality context conflict evaluation setting, where an MLLM is asked to answer a question based on a pair of contrasting evidence.
+Multimodal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video).
+Recent advances demonstrate their exceptional capabilities on complex tasks with multimodal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multimodal dialogue systems (Sun et al., 2022).
+Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multimodal inputs.
+To investigate this, one line of work (Fu et al., 2024; Amara et al., 2024) compares model performance on unimodal input, providing either only text or only image input for the same question.
+Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes of the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu & Frank, 2024; 2022).
+However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multimodal scenarios, where information from different modalities naturally co-occur.
+In this paper, we provide a controllable setup to study the modality preference in MLLMs.
+As shown in the left panel of Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting evidence from different modalities.
 In this way, we can determine the modality preference based on the answer given by MLLMs.
-Moreover, to dismiss the effect of several confounding factors, including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we utilize widely adopted images and select samples that demonstrate accurate question comprehension and reliable single-modality perception.
-Building upon this, we introduce a semi-automated annotation framework to construct the refined Modality Context Conflict dataset, MC2, which covers eight recognition-level tasks with 2,000 carefully selected samples.
-Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 18 representative MLLMs.
+To dismiss the effect of confounding factors including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we annotate and select perception-level tasks that demonstrate accurate question comprehension and reliable single-modality recognition.
+Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight perception-level tasks with 2,000 carefully selected samples.
+Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 20 representative MLLMs.
 Our study reveals several intriguing findings:
-  • All tested MLLMs generally demonstrate clear modality bias, as illustrated in the right penal of Figure 1, and the modality bias is affected by model-intrinsic factors and task-specific characteristics.
-  • The direction of modality preference in MLLMs can be modulated through external interventions, such as modality degradation or instructional guidance.
-  • Latent space analysis confirms that modality preferences manifest as geometrically separable patterns in MLLMs’ representations, which can be captured by distinct vectors.
-Built on these, we propose a modality preference probing and steering method based on representation engineering to explicitly amplify the direction of modality preference without the need for additional fine-tuning or prompt design.
-Experimental results show that the proposed method leads to notable performance improvements on both hallucination mitigation and multimodal machine translation.
+  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the right panel of Figure 1, and modality preference can serve as a useful indicator of downstream task performances.
+  • Internal attention patterns toward specific modalities give rise to modality preference, and the underlying factors can be traced to the training data recipe and the scale of the MLLMs.
+  • Modality preference can be modulated through explicit instruction guidance, and the direction of preference can be captured as geometrically separable patterns in the latent space.
+Built on these, we propose a modality preference probing and steering method based on representation engineering (Zou et al., 2023) to explicitly amplify the modality preference without additional fine-tuning.
+Experimental results show that the proposed method leads to notable performance improvements on multimodal visual understanding and multimodal machine translation.
 Our main contributions are summarized as follows:
-  • We introduce a Modality Context Conflict benchmark for the systematic evaluation of modality preference in MLLMs.
-  • Our systematic analysis reveals intrinsic modality preferences in MLLMs, and we find that such preferences are steerable and can be identified through latent representations, offering insights into reasoning under multimodal conflicts.
-  • Inspired by these findings, we propose a method to steer modality preference via representation-level intervention, enabling controllable preference adjustment on MC2 and improving performance on downstream tasks.
 Figure 1: Illustrations of evaluating modality preference.
-Left: Use multimodal conflict pairs to evaluate modality preference.
-Right: Quantified results for modality preferences.
-Vision and Text represent vision and text preference, other represents irrelevant prediction.
+Left: Using multimodal conflict context pairs to evaluate modality preference.
+Right: Quantified scores for Vision and Text modalities, where a higher score indicates a stronger preference toward the corresponding modality.
+Other represents irrelevant predictions, which are discarded during the statistics.
+  • We introduce MC2 to comprehensively evaluate modality preferences in MLLMs and highlight the significance of modality preference in correlating the downstream task performance.
+  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representation, providing insights into multimodal reasoning.
+  • We propose a training-free method that steers modality preference via representation-level intervention, enabling controllable preference adjustment and enhancing performance on downstream tasks.
 2 Related Work
-Multimodal Large language Models.
-In recent years, Large Language Models (LLMs) have achieved impressive results across a wide spectrum of tasks [33, 53, 15, 47, 19].
-Building upon these, Multimodal Large Language Models (MLLMs) have extended such capabilities, demonstrating strong performance in visual question answering, structured information understanding, and visual navigation [55, 51, 32, 44, 52].
-Recent research has turned to systematically evaluating their biases and failure modes [30, 1, 5, 4, 9], aiming to better understand their limitations and guide the development of more reliable models.
-Modality Preference in MLLMs.
-Modality preference—whether MLLMs prioritize one modality over another and to what extent they favor a specific modality when resolving multimodal inputs—can reveal how MLLMs jointly reason over different modalities.
-Some studies have also observed that MLLMs often favor textual inputs while neglecting visual information, which can result in hallucinations [8, 34, 54].
-This highlights the importance of investigating modality preference.
-To study the modality preferences of MLLMs, one common approach involves evaluating model performance on unimodal inputs—comparing results when only text or only image inputs are provided for the same task [12].
-However, this method inherently overlooks how models handle genuinely multimodal scenarios where both modalities are presented simultaneously, thus limiting its effectiveness in capturing true modality preference.
-Another research direction assesses the relative influence of textual and visual inputs using Shapley value-based attribution methods [4, 36, 37].
-Yet, we argue that both modalities are typically essential for solving multimodal tasks, making this approach insufficient for identifying a model’s inherent modality preference in realistic multimodal contexts.
-Concurrent work evaluates the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts [45].
-However, conflict detection is only one aspect of multimodal reasoning and does not fully reflect a model’s overall ability to jointly reason across modalities.
-In this work, in order to investigate their modality preferences, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts.
-Compared to prior work, we carefully control for potential confounding factors that may influence modality preference, including question understanding, unimodal content recognition, and interference from the model’s internal knowledge.
-We systematically evaluate modality preference and design a method to controllably steer it.
-Finally, we demonstrate the effectiveness of our approach on downstream tasks.
-Representation Engineering and Activation Editing.
-Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in activation space [31, 35, 39, 42].
-Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs [14, 38, 46, 48, 56].
+2.1 Modality Preference
+Existing studies on the modality preference of MLLMs can be broadly divided into two categories: 1) investigating the data-related factors that give rise to modality preference or bias, and 2) analyzing the intrinsic characteristics of modality preference within models.
+Data factors influencing modality preference.
+Research on data-related factors (Guo et al., 2023; Chen et al., 2024; Leng et al., 2024) explores how properties of multimodal datasets give rise to and reinforce modality preference.
+In particular, many samples in multimodal datasets can be resolved correctly by relying on information from only a single modality.
+When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024).
+Furthermore, their inclusion in evaluation benchmarks artificially inflates performance metrics, as models can exploit these unimodal shortcuts instead of performing genuine cross-modal integration (Leng et al., 2024; Winterbottom et al., 2020).
+While these studies establish data’s role in inducing preference, our work focuses on exploring the intrinsic modality preference inherent in MLLMs themselves, independent of specific data distributions.
+Evaluating the intrinsic modality preference in MLLMs.
+Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multimodal models optimize for multimodal inputs.
+Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024).
+While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multimodal systems.
+Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu & Frank, 2022; 2024).
+However, in real-world multimodal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining truly modality preference.
+Wu et al.
+(2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts.
+However, conflict detection is only one facet of multimodal reasoning and does not comprehensively reflect a model’s modality preference when processing multimodal contexts.
+In this work, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts.
+Compared to prior work, we carefully control confounding variables such as input quality, question-understanding ability, and internal model knowledge, and construct a modality context conflict dataset, enabling a more rigorous evaluation of modality preference and uncovering new insights.
+Furthermore, we design a flexible method which can controllably steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
+2.2 Representation Engineering
+Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023).
+Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs (Greenblatt et al., 2023; Stolfo et al., 2024; Wu et al., 2024; Xu et al., 2024; Zou et al., 2023).
 In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs).
-Instead of focusing on abstract properties, our method identifies and manipulates activation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
+Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
 3 The MC2 Benchmark
 In this section, we introduce the design and methodology behind the construction of the Multimodal Context Conflict dataset, MC2, intended for evaluating modality preference.
-We outline the data design philosophy in Section 3.1, followed by the data construction pipeline in Section 3.2 and the evaluation method in Section 3.3.
+We outline the data design philosophy in Section 3.1, followed by the data construction pipeline in Section 3.2 and the question design and evaluation metric in Section 3.3.
 3.1 Data Design Philosophy
-To evaluate modality context preference, we construct modality context conflict pairs that simulate how contexts from different modalities influence model decisions, thereby revealing modality preference.
-However, the evaluation process may be influenced by various factors such as the internal knowledge within MLLMs, question comprehension and the ability to understand single-modality context.
-Therefore, on the one hand, we focus on evaluating recognition-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs.
-On the other hand, we ensure that the evaluated models can accurately answer questions based solely on the context of a single modality, minimizing the influence of confounding factors.
-Definitions and Notations: The following definitions and notations are used throughout this paper.
-In this paper, the image is selected as the vision context,$c^{v}$.
-We utilize advanced LLMs to generate textual contexts $c^{t}$ that conflict with the vision context $c^{v}$ in relation to the question $q$.
-$A^{v}$ and $A^{t}$ are answers based on the vision context or textual context.
+To reveal the modality preference of MLLMs when processing truly multimodal inputs, we construct modality context conflict pairs and carefully control various confounding factors that may affect the measurement of modality preference, such as the internal knowledge of MLLMs, question comprehension, and the understanding of unimodal contexts.
+Therefore, we focus on evaluating perception-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs.
+Besides, we ensure that the evaluated models can accurately answer questions based solely on the context from a single modality, minimizing the influence of confounding factors.
 3.2 Semi-automated Data Construction Pipeline
-In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design principles outlined in Section 3.1.
-The dataset is derived from the TDIUC [22] dataset, sourced from MS-COCO [26], widely adopted in model development to ensure the evaluated models can recognize the images.
-We extract image caption $caption$, vision context $c^{v}$, question $q$, and answer $A^{v}$ for each sample from TDIUC as the foundation for data annotation.
+In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design philosophy outlined in Section 3.1.
+The dataset is derived from the TDIUC (Kafle & Kanan, 2017) dataset, sourced from MS-COCO (Lin et al., 2014), widely adopted in model development to ensure the evaluated models can recognize the images.
+We select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the vision context and the image caption capcap for each sample from TDIUC as the foundation for data annotation.
 The pipeline follows these steps:
-Candidate Textual Context Generation.
-We use DeepSeekV3 [28] and ChatGPT4o-mini [20] to generate distractor answers $A^{t}$ and candidate textual contexts $c^{t}$ supporting the distractor answers.
-We aim to construct textual contexts that are aligned with the visual context in terms of overall scene semantics, while introducing content-level conflicts with the visual context specifically in relation to the given question.
-Therefore, the following task template is employed to generate the distractor answers and textual contexts:
-USER: <caption> <question> <answer based on vision context>.
-Generate the distractor answer different from the given answer based on the vision context and generate a textual context to support the distractor answer.
-Ensure that the scene described in the textual context is consistent with the caption.
-ASSISTANT: <output>
-For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection and construction. Detailed prompt templates for different LLMs are provided in Appendix B.1.
-Context-based Sample Selecting.
-We use several basic judge MLLMs, such as LLaVA1.5-7B [29] and QwenVL-7B [6], to select samples.
-This ensures that all subsequent MLLMs to be evaluated demonstrate high recognition capabilities for both vision and textual contexts.
-The following task template is used to prompt the judge MLLMs for selecting samples:
-USER: <vision/textual context> <question>. Answer question based on the context.
-ASSISTANT: <output>
-We select only those samples for which all judge MLLMs can correctly answer the given question based on either the vision context or the textual context
+Textual Context Construction.
+Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with the $c^{v}$ and capcap in terms of overall scene semantics.
+We prompt DeepSeekV3 (Liu et al., 2024a) and ChatGPT4o-mini (Hurst et al., 2024b) to generate a distractor answer $A^{t}$ to $q$, together with $c^{t}$ that plausibly supports $A^{t}$, using carefully crafted instructions.
+For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection.
+To ensure that all evaluated MLLMs demonstrate strong recognition capabilities for both visual and textual contexts, we employ several basic MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
 Human Verification.
-Although the previous steps generally yield reliable results, we incorporate manual inspection to ensure the high quality of the data annotation.
-In this step, we verify the existence of conflicts between $c^{v}$ and $c^{t}$ and ensure that both contexts can correctly direct $q$ to the corresponding answers, $A^{v}$ and $A^{t}$.
-When errors are found in a sample, we either correct or discard the sample entirely.
-Each category is cross-verified by three annotators to ensure the reliability of the results.
-The detailed manual annotation tasks and implementation procedures are provided in Appendix  B.1.
+We incorporate manual inspection to ensure the high quality of the data annotation.
+Specifically, we verify the existence of conflicts between $c^{v}$ and $c^{t}$ and ensure that both contexts can correctly direct $q$ to the corresponding answers, $A^{v}$ and $A^{t}$.
+Each sample is cross-verified by three human annotators to ensure the reliability of the results, and when errors are found, annotators either correct or discard the sample entirely.
 Iterative Refinement.
-The dataset undergoes multiple rounds of refinement through a feedback loop.
-This process ensures that the conflicts between different modality contexts are preserved and consistently represented, and MLLMs can correctly understand both the vision and textual contexts.
-The iterative process helps identify and rectify potential errors, thereby enhancing the dataset quality.
-To this end, the dataset is organized into eight categories: positional reasoning, counting and color, attribute recognition, sentiment analysis, activity recognition, sports recognition, and object recognition, with 250 samples in each category.
-Details of the data annotation format along with sample cases and dataset statistics is presented in Appendix B.3.
-3.3 Question Design and Preference evaluation
+The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which helps identify and rectify potential errors, thereby enhancing the dataset quality.
+Modality Context Conflict Dataset. To this end, we construct MC2, a modality context conflict dataset including 2000 samples.
+The detailed instruction templates for textual context generation, the detailed manual annotation procedures, the data annotation format along with sample cases and dataset statistics are provided in Appendix B.
+3.3 Question Design and Evaluation Metric
 Question Design.
-We reformulate the original questions using GPT-4o-mini [21] into binary-choice format.
-Additionally, to mitigate potential position bias in model outputs, we implement a consistency evaluation strategy.
-Specifically, for each binary-choice question, we create two versions by swapping the order of the answer choices such as, “How many people are there?
-A.
-four.
-B.
-five.” and “How many people are there?
-A.
-five.
-B.
-four.” A model’s response is considered consistent only if it selects different answers in the two versions (i.e., A then B, or B then A).
-If the model selects the same option in both versions, the response is marked as invalid, indicating possible position bias.
-Evaluation for modality preference.
-To quantify modality preference, we evaluate whether the model’s prediction is based on visual or textual context according to the following three types:
-  • Vision: The prediction aligns with visual context, indicating a preference for vision modality.
-  • Text: The prediction aligns with textual context, indicating a preference for text modality.
-  • Others: The model gives inconsistent responses or outputs irrelevant answers. These instances are discarded from further analysis.
-We use the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/{S_{\text{vision}}+S_{\text{text}}}$.
-Vision Ratio greater than 0.5 indicates that the model tends to favor visual information over text.
-Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-LLaMAVision 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
-LLaVA1.5-13B 34.4/59.6 8.8/89.6 4.8/88.0 12.0/84.8 1.6/96.4 12.8/82.0 9.6/87.6 31.2/62.8 14.4/81.3
-OneVision-7B 32.0/36.4 21.6/54.8 2.8/94.4 24.8/56.4 2.4/86.4 30.0/38.0 11.6/71.2 42.4/31.2 20.9/58.6
-Owl3-7B 60.8/31.6 16.4/72.4 10.8/85.6 22.0/69.6 8.4/88.0 28.4/60.4 17.2/71.2 60.0/29.6 28.0/63.5
-Qwen2VL-7B 26.4/58.0 12.4/82.8 0.8/95.6 13.2/80.4 4.0/93.6 16.0/78.8 11.6/83.6 38.0/54.0 15.3/78.3
-Qwen2.5VL-7B 65.6/12.8 45.2/46.0 18.0/68.8 46.4/38.0 51.6/39.6 70.8/20.0 42.0/43.6 77.6/14.0 52.2/35.4
-GLM-4V-9B 42.0/42.4 32.4/59.2 8.8/81.6 28.0/62.4 15.2/74.4 56.8/32.8 23.3/66.0 54.0/32.8 32.6/56.5
-SPHINX-V2-1K 39.6/50.8 14.8/82.4 1.2/98.4 16.8/77.6 9.2/85.6 23.2/69.2 24.4/67.2 59.2/32.4 23.6/70.5
-InternVL3-9B 45.2/35.2 21.2/68.0 20.8/62.4 27.2/54.4 23.2/50.4 38.0/40.4 19.6/63.2 76.8/14.8 34.0/48.6
-InternVL3-14B 72.8/8.8 30.8/48.4 25.2/60.0 33.2/52.0 37.2/47.2 58.0/21.2 24.8/52.8 84.4/9.6 45.8/37.5
-CogVLM2-19B 44.0/39.6 29.2/56.0 8.8/75.6 19.2/54.8 8.0/73.2 31.6/43.2 25.2/60.8 59.2/28.4 28.2/54.0
-LLaVA-next-7B 10.8/74.4 5.2/85.2 0.8/93.2 3.6/79.6 0.4/90.8 6.0/76.0 4.8/73.6 26.0/46.8 7.2/77.5
-GPT-4o-mini 94.4/3.2 35.6/47.6 60.4/28.4 22.0/58.9 19.4/59.2 34.8/36.4 71.2/20.4 78.4/12.8 52.0/33.4
-Table 1: Results for modality preference. Values are reported as vision/text results for each model.
-4 What Modality Preference Has MLLMs Developed?
-We use the proposed MC2 dataset to evaluate a wide range of MLLMs, including the proprietary GPT-4o-mini [21] and seventeen open-source MLLMs with varying parameter sizes.
-For detailed model evaluation, please refer to Appendix C.
-We investigate the modality preferences of MLLMs by addressing two key questions: Which modality do MLLMs prioritize?
-and Can modality preference be controlled?
-We first analyze the inherent modality preferences of different MLLMs, and then delve into how such preferences can be controlled.
+We reformulate the original questions with ChatGPT-4o-mini (Hurst et al., 2024a) into a binary-choice format.
+To further reduce potential position bias in multimodal inputs, we adopt a consistent evaluation strategy, similar to Liu et al.
+(2024d).
+Concretely, for each question, we construct two versions by swapping the order of the answer choices.
+A model’s prediction is regarded as consistent only if it selects the same answer in both versions for a sample; otherwise, it is labeled as inconsistent.
+Such inconsistent samples are discarded from the subsequent measurement of modality preference.
+Evaluation Metric.
+Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this idea to evaluate the modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities.
+More importantly, through the careful design of our benchmark, we establish as a basis that the model can reliably understand both modalities in isolation.
+As shown in Table 9 and Table 10 in Appendix, all models achieve over 95% accuracy when provided with either textual or visual context.
+Building on this, our metric evaluates modality preference by assessing how the model’s responses align with textual or visual input when the two provide conflicting signals.
+The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the responses are ambiguous, uncertain, or inconsistent with either modality, which are discarded from further analysis.
+Then, we naturally define the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/({S_{\text{vision}}+S_{\text{text}}})$, where $S_{\text{vision}}$ or $S_{\text{text}}$ denotes the score of the vision or text modality, computed as the proportion of samples whose responses are categorized as Vision or Text across the dataset.
+$S_{others}$ is the proportion of samples whose responses are categorized as Others.
+Vision Ratio greater than 0.5 indicates that the model tends to favor visual context over text.
+4 Modality Preferences in MLLMs
+This section presents a systematic investigation into modality preference in MLLMs, structured around four key research questions: 1) Which modality do MLLMs prioritize?
+2) What factors drive these preferences?
+3) Can the Vision Ratio provide guidance for downstream task performance?
+4) Can modality preference be controlled?
 This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
+Figure 2: Results of modality preference across different MLLMs.
+Left: Quantified scores for Vision, Others Text modalities using $S_{vision}$, $S_{others}$ and $S_{text}$ as well as Vision Ratio.
+Right: Trends of Vision Ratio with respect to model parameter size across different MLLMs.
 4.1 Which Modality Do MLLMs Prioritize?
-Before evaluating modality preference, we first assess the ability of MLLMs to answer questions accurately given a single-modality context in the MC2 dataset.
-All models achieve over 95% accuracy when provided with either textual or visual context.
-Please refer to Appendix C.1 for detailed results.
-Figure 2: Results of modality preference. Left: averaged modality preferences of different MLLMs, Right: vision Ratios for different MLLMs across various tasks.
-How do different MLLMs exhibit preferences across various tasks?
-As described in Section 3.3, we quantify modality preference using model predictions categorized as Vision, Text, and Others, and use the Vision Ratio to quantify the preference for the vision modality, with the results presented in Figure 2 and Table 1.
-We observe that all MLLMs demonstrate some level of modality preference bias.
-We observe that the majority of MLLMs exhibit a strong preference for the text modality, particularly as evidenced by the Vision Ratio of LLaVA1.5-13B, which is only 14.4% in the left panel of Figure 2.
-Interestingly, the Qwen2.5VL and InternVL3 models show a certain degree of preference for the vision modality, which contrasts with the previous conclusion that MLLMs suffer from a severe language prior [23, 37, 45].
-A more detailed analysis of different tasks reveals that Qwen2.5VL and InternVL3 have higher Vision Ratios across all tasks compared to other models, especially in high-level recognition tasks such as sports and object recognition, as shown in the right panel of Figure 2.
-Moreover, we find that low-level tasks (e.g., attribute and color recognition) as well as tasks involving more complex reasoning (such as sentiment analysis, positional reasoning, and counting) consistently exhibit a stronger reliance on the text modality across all MLLMs.
-What is the relationship between modality preference and model size?
+We use the MC2 benchmark to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
+Different MLLMs exhibit different modality preferences.
+As described in Section 3.3, we quantify modality preference using Vision Ratio, with the results presented in the left panel of Figure 2 and detailed in Table 6.
+We observe that all MLLMs exhibit clear modality preference, with most models showing a strong preference for text; for instance, LLaVA1.5-7B attains only a 13.4% Vision Ratio.
+This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu & Frank, 2024; Wu et al., 2025).
+Interestingly, the Qwen2.5VL and InternVL3 show a certain degree of preference towards the vision modality.
+Larger MLLMs exhibit stronger preferences for the vision modality.
 We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference.
-As shown in the left panel of Figure 3, we observe that for all model families, the preference for the vision modality increases with the model size.
+As shown in the right panel of Figure 2, we observe that for all model families, the preference for the vision modality increases with the model size.
 And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases.
-However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their size increases.
-Figure 3: Evaluation results of modality preference and attention analyse.
-The left shows modality preference of MLLMs as parametersize changes, measured by Vision Ratio.
-A higher Vision Ratio indicates a stronger vision modality preference.
-the right shows the trends of Vision Ratio and multimodal Attention Ratio across different tasks and models.
+However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their sizes increase.
+Vision Ratio aligns with human preference.
+We further verify whether the Vision Ratio can serve as a human-level measure of modality preference in MLLMs.
+We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, LLaVA-OneVision-7B, InternVL3-14B, and LLaVA1.5-7B.
+In addition, we craft prompts to elicit explicit reasoning chains from the models, specifically targeting their reliance on visual or textual information.
+To ensure labeling reliability, three expert annotators independently annotate the expressed modality preference for each response, with the final label determined by majority vote.
+The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with the ones given by human ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%.
+This indicates that the Vision Ratio can act as a reliable, automated proxy for human assessment of modality preference.
+Figure 3: Analysis of modality preferences.
+Left: Trends of Vision Ratio and multimodal Attention Ratio across different models.
+Middle: Vision Ratio with Respect to the Proportion of Multimodal Conflict-Context Training Data and different MLLMs.
+Right: Relationship between visual understanding ability, quantified as the average accuracy across seven widely used benchmarks and the modality preference measured by Vision Ratio.
+4.2 What Factors Drive These Preferences?
+Given that different MLLMs display varying modality preferences, we examine two primary sources of such differences: the internal attention distribution and the training factors.
+Different allocation of attention across modalities.
+We compute the mean attention scores over all token positions from both modalities and define the ratio of visual attention to total attention as the Attention Ratio.
+By analyzing Qwen2.5VL-7B and LLaVA-OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3.
+This alignment suggests that MLLMs distribute attention unevenly between modalities, which in turn contributes to their divergent modality preferences.
+Impact of model scale and training data recipe.
+Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM.
+Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multimodal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference.
+2) Larger LLMs are more capable of shifting their preference during training;
+To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multimodal conflict contexts, adjusting their preferences toward text or vision.
+We then measured changes with the Vision Ratio.
+We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio.
+As shown in the middle panel of Figure 3, increasing the proportion of multimodal contexts consistently leads to larger preference shifts, supporting Hypothesis 1.
+This suggests that multimodal inputs maybe create more challenging training conditions, leading to stronger shifts of preference.
+Furthermore, Qwen2.5VL-7B exhibites greater shifts than Qwen2.5VL-3B under the same conditions, supporting Hypothesis 2.
+This indicates that larger LLMs demonstrate stronger learning ability and adapt more effectively.
+4.3 Can the Vision Ratio provide guidance for downstream task performance?
+To demonstrate the correlation between the modality preference and performance for downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs.
+Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in the Appendix C.2.
+We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3.
+The results reveal a strong positive association between the Vision Ratio and visual understanding ability across the evaluated MLLMs.
+Specifically, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that the Vision Ratio provides a highly reliable indicator of visual understanding task performance.
 4.4 Can Modality Preference Be Controlled?
-We employ two intervention methods, content quality manipulation and instruction-based guidance, to explore whether modality preference can be controlled.
-Figure 4: Effects of different interventions on the modality preference, measured by Vision Ratio.
-The left two panels: modifying input quality, including image noise and text grammar errors.
-The right two panels: using specific instructions to guide MLLMs toward the vision or text modality.
-Further analysis using attention attribution.
-To further investigate the internal mechanisms underlying modality preference, we analyze the attention distributions over the vision and textual contexts in Qwen2.5VL-7B and LLaVA-OneVision-7B.
-Specifically, we compute the average attention scores for each modality at every token position across all transformer decoder layers of the MLLMs, and define the ratio between visual and textual attention as the Attention Ratio.
-We then visualize the Attention Ratio alongside the Vision Ratio across a set of representative tasks.
-As shown in right panel of Figure 3, the trends of Vision Ratio and Attention Ratio are highly aligned across both models and tasks.
-This suggests that modality preferences are closely related to the internal attention allocation.
-A similar observation has also been reported in prior work [45].
-How does modality content quality affect modality preference?
-We manipulate the quality of input modality content by introducing Gaussian noise to visual contexts and adding grammatical or spelling errors to textual contexts to observe shifts in modality preference.
-As shown in the left two panels of Figure 4, we observe that adding noise to images causes a shift in modality preference toward the text modality, whereas introducing grammatical or spelling errors into the text leads to a shift toward the vision modality.
-It is noteworthy that, despite the degradation in the quality of one modality, the models’ ability to comprehend information from a single modality remains largely unaffected.
-This implies that MLLMs may implicitly associate modality reliability with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning.
-Details of the data generation, experimental design, and full results are provided in Appendix C.2.
-Figure 5: PCA visualization of modality preference directions under different intervention strategies.
-Left: representation shifts under noise-based interventions (image noise or textual errors).
+We employ instruction guidance to investigate whether modality preference can be controlled, and conduct a latent space representation analysis to examine the mechanisms, underlying the preference adjustment.
+Details of the experimental design and results are provided in Appendix C.3.
+Figure 4: Analysis of modality preference under instruction-guidance.
+Left: Adjustment of modality preference using instruction-guided control (Inst-vision vs.
+Inst-text).
 Middle: Representation shifts under instruction-guided interventions.
-Right: layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction.
-Can modality preference be guided through instruction design? We further investigate the impact of instruction design on modality preference. We design the following task instruction:
-USER: <vision context> <textual context> <question>.
-In this case, there is a conflict between the vision and textual context, and you should rely on the textual/(vision) context rather than the vision/(textual) context.
-ASSISTANT: <output>
-As shown in the right two panels of Figure 4, instructions directing the model toward either the textual or vision modality effectively guide the modality preference in the specified direction.
-Moreover, guidance toward the textual modality tends to produce a more pronounced effect compared to guidance toward the vision modality.
-This tendency may reflect the disproportionate exposure to textual instructions during training, leading to stronger alignment with text-guided cues.
-Modality Preference Direction in Representation Space To further understand how the two intervention methods influence modality preference internally, we analyze the hidden representations of the models.
-Specifically, we apply Principal Component Analysis (PCA) to the hidden states across different layers to identify the dominant direction corresponding to modality preference shifts.
-Detailed implementation is described in Appendix C.3.
-As shown in the left panel of Figure 5, introducing Gaussian noise to visual inputs leads to a noticeable shift in the representation space along a clear direction, while adding grammatical or spelling errors to text inputs results in a comparatively weaker shift.
-This asymmetry is consistent with our earlier observations in Section 4.2, where textual perturbations steer less pronounced changes in modality preference compared to visual perturbations.
-Furthermore, the middle panel of Figure 5 illustrates that instruction-based interventions also steer significant representation shifts, aligning with the modality specified in the instruction.
-Notably, the directionality in the PCA space reveals that the model’s internal states are indeed sensitive to modality control cues, whether they originate from instruction design.
-Inspired by this, in the following section 5, we propose a probing and Steering method to amplify modality preference and apply it to downstream tasks.
+Right: layer-wise absolute difference and standard deviation of hidden states between different instruction.
+Modality preference can be guided through instruction design.
+We investigate the impact of instruction design on modality preference, for instance, by explicitly directing the model to rely on a specific modality when answering a question.
+Specifically, we evaluate the modality preference measured by Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text or vision preference instruction (Inst-text, Inst-vision).
+As illustrated in the left panel of Figure 4, instructions that steer the model toward a particular modality effectively shape its modality preference.
+Modality preference direction in representation space.
+To further understand how the intervention methods influence modality preference internally, we analyze the hidden representations of the models.
+Specifically, we apply Principal Component Analysis (PCA; Abdi & Williams, 2010) to the hidden states to identify the dominant direction corresponding to modality preference shifts.
+The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, aligning with the modality specified by the instruction.
+The PCA direction further reveals that the model’s internal states are sensitive to modality control cues, which motivates us to develop representation techniques for adjusting modality preference and to apply these insights to downstream tasks in Section 5.
 Figure 5: Overall framework of the proposed method.
-Modality Preference Probing collects the neural activity representations in response to each modality preference prompt.
-It then computes and scales the direction of modality preference that expresses a preference for a specific modality.
+Modality Preference Probing collects the neural activity, computes and scales the direction of modality preference.
 Modality Preference Steering selects the target layer during the second inference and adds the scaled modality preference direction to the representation at the corresponding layer at each inference step.
-5 Representation-Based Modality Steering
-In this section, we adapt the concept of representation engineering [56] to probe the modality preference direction and reveal the consistent representation behavior behind both interventions discussed in Section 4.2.
-Inspired by this, we steer and amplify the preference direction to control the model’s behavioral expression.
-Representation engineering involves collecting neural activity by constructing vectors of activation values that cause desired changes to output text, and controlling the model’s behavior by adding the probing vectors to the forward passes of a frozen LLM.
-As shown in Figure 6, the proposed framework consists of two parts: Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
+5 Representation Based Modality Preference Steering
+Inspired by the representation behavior discussed in Section 4.4, we propose to use the representation engineering (Zou et al., 2023) to steer the modality preference, controlling the model’s behavioral expression.
+As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
 5.1 Modality Preference Probing
-Modality Preference Probing (MPP) is proposed to probe and collect neural activity that represents the direction of modality preference.
-Inspired by the pre-training objective of decoder-only MLLMs, Next Token Prediction objective [21] and previous works that extract classification features from the last token [11], we collect neural activity from the last token in the input text.
-The process involves probing modality preference through two pairs of modality preference requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context rather than the text context’) and another with a text preference probing (e.g., ‘answer the question based on the text context rather than the vision context’).
+We probe and collect neural activity that represents the direction of modality preference.
+Inspired by the pre-training Next Token Prediction objective of decoder-only MLLMs (Hurst et al., 2024a) and the method to extract classification features (Feucht et al., 2024), we collect neural activity from the last token in the input text.
+The process involves probing modality preference through two pairs of requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’).
 Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$.
@@
 We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs.
-More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards expressing the text modality at layer $l$ for a given query as:
+More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $l$ for a given query as:
 $\mathbf{u}^{t}_{\ell}=\frac{1}{N}\sum_{i}^{N}\big(\mathbf{x}_{i,\ell}^{t}-\mathbf{x}_{i,\ell}^{v}\big).$ (1)
-Averaging over different queries allows us to capture the activation values most closely associated with modality preference requests, independent of questions.
+Averaging over different queries allows us to capture the activation values most closely associated with modality preference, independent of questions.
 As shown in the right panel of Figure 4, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples.
 We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers.
-Similar patterns are observed for Qwen2VL and LLaVA-OneVision, as detailed in Appendix D.1.
-Based on this observation, we select the corresponding layers of the model to control the direction of modality preference in Section 5.2.
-Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-Vision MLLM-only 26.4 12.4 0.8 13.2 4.0 16.0 11.6 38.0 15.3
-Inst Design 60.8 24.0 20.0 20.4 10.8 32.0 27.2 63.2 32.3
-CoT Prompting 57.6 27.6 16.0 18.0 21.6 35.6 44.4 52.4 34.2
-Few shot 32.0 12.0 10.0 11.0 2.0 25.0 9.0 38.0 17.2
-ours 78.8 35.6 38.8 29.6 22.4 56.4 45.2 78.4 48.1
-Text MLLM-only 12.8 46.0 68.8 38.0 39.6 20.0 43.6 14.0 35.4
-Inst Design 14.4 46.8 72.8 40.4 55.6 24.4 35.6 11.6 37.7
-CoT Prompting 27.2 63.6 83.6 62.8 75.6 53.2 58.0 20.4 55.6
-Few shot 21.0 77.0 89.0 73.0 73.0 60.0 70.0 42.0 63.1
-ours 69.6 67.6 84.4 50.8 82.8 57.6 54.8 41.2 63.6
-Table 2: Performance by steering Qwen2VL-7B towards vision and Qwen2.5VL-7B towards text.
+Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 5.2.
+Similar patterns are observed for Qwen2VL-7B, Qwen2.5VL-7B, LLaVA-OneVision and InternVL3, as detailed in Appendix D.1.
 5.2 Modality Preference Steering
-After probing the direction of the modality preference, we compute the steering vector by re-scaling the direction vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$.
-The scaling process must carefully balance two objectives: it must be strong enough to effectively steer the model’s modality preference, while simultaneously preserving the model’s normal output behavior.
-In our preliminary experiments, we observe that setting the weight too large led to repetitive and meaningless outputs, whereas a weight that is too small fails to steer any noticeable change in the model’s modality preference.
-Unlike previous approaches [56, 38] that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original hidden state distribution.
+After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$.
+The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior.
+In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference.
+Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states.
 This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities.
@@
 $\mathbf{s}^{t}_{\ell}=w\mathbf{u}^{t}_{\ell},\ \text{where}\ w=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\mathbf{x}^{t}_{i,\ell}\|}{\|\mathbf{u}^{t}_{\ell}\|}$ (2)
-Finally, during a new round of inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ (as identified in Section 5.1) at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens—including vision tokens, text tokens, and newly generated tokens—to steer the model’s response toward the text modality.
-Similarly, to steer responses towards the vision modality, the adjustment is performed in the opposite direction.
-In the final implementation of our method, no additional data or labels are introduced during either the probing or steering stages.
-We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the model’s modality preference.
-Notably, the use of unlabeled or self-generated signals avoids annotation bias, which is a critical property when aiming to enhance the generalization ability of different models across diverse domains.
-5.3 Verifying the Effectiveness of Steering Modality Preference
-We verify the effectiveness of the proposed method in controlling modality preference on MC2 across several MLLMs, including Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-14B.
-We consider several training-free approaches as baselines: MLLM-only refers to MLLM directly reasoning in modality-conflicting contexts; Inst-Design uses instructions to guide modality preference direction; CoT-prompting enables complex reasoning through intermediate steps; and Few-shot uses a few examples to guide the model in answering.
-For detailed method implementation and the results of LLaVA-OneVision-7B and InternVL3-14B, refer to Appendix D.2.
-As shown in Table 2, the proposed method significantly outperforms the baseline approaches, particularly in high-level recognition tasks.
-Phd Model Attribute Sentiment Positional Counting Object Avg
-Phd-icc MLLM-only 10.0 2.50 3.50 6.0 8.0 6.0
-Inst Design 14.5 2.5 1.5 5.5 25.0 9.8
-CoT Prompting 5.0 6.0 8.50 6.5 40.5 13.3
-Few shot 3.0 0.5 1.5 5.0 2.0 2.4
-+ours 10.0 11.0 14.0 5.0 51.5 18.4
-Phd-iac MLLM-only 28.5 8.5 20.5 30.5 50.0 27.6
-Inst Design 34.5 13.0 26.0 39.0 60.0 34.5
-CoT Prompting 15.5 23.5 30.2 17.0 59.0 29.0
-Few shot 17.0 9.0 14.5 29.0 37.0 21.3
-+ours 29.5 14.5 35.0 28.5 63.5 34.2
-Table 3: Performance of the proposed method on Qwen2VL-7B using the hallucination evaluation dataset from Phd [30].
-Phd-icc provides an incorrect context in addition to the correct image, while Phd-iac introduces an inaccurate context to assess model robustness under misleading input.
+Finally, during the second inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward the text modality.
+Similarly, steering models towards vision modality is performed towards the opposite direction.
+In the final implementation, no additional data or labels are introduced.
+We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the modality preference.
+Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality and steering Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{vision}$ and $S_{text}$.
+Preference Model MLLM InstDesign CoT FewShot Ours
+[…]
+OneVision-7B 20.9 31.6 36.7 38.2 42.8
+[…]
+InternVL3-8B 37.5 52.8 53.1 49.8 57.1
+Table 2: Multimodal translation results for Ambigcaps (Li et al., 2021). BLEU scores are reported for English (En) $\leftrightarrow$ Turkish (Tr).
 Method En->Tr Tr->En
@@
       +Inst towards text  9.45 (+0.53) 18.98 (+0.42)
-      +ours 10.22 (+1.3) 19.89 (+1.33)
-Table 4: The translation results for Ambigcaps using different instruction for preference towards vision modality and text modality and the proposed method steering modality preference towards text modality.
-We report the Bleu score for the bidirectional translation between English (en) and Turkish (tr).
-5.3.1 Downstream Task Applications of Modality Preference Control
-We also evaluate the effects of adjusting modality preference towards specific directions on two downstream tasks.
-Specifically, we assess the hallucination mitigation effect on two subsets of the PhD dataset [30], namely Phd-icc and Phd-iac.
-Additionally, we evaluate the performance on a multimodal machine translation (MMT) task using the AmbigCaps [25] dataset.
-By steering modality preference towards the vision modality, which places greater trust in visual information, we mitigate multimodal hallucinations.
-As shown in Table 3, our method significantly outperforms the baseline methods on PhD.
-Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation.
-This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output.
-We use our method to steer the modality preference towards the text modality.
-As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish.
-For detailed implementation and results, refer to Appendix D.3.
+      +Ours 10.22 (+1.30) 19.89 (+1.33)
+Table 3: Performance of the proposed method on the visual understanding benchmark, Phd (Liu et al., 2024c). we report the accuracy results on the phd-icc/phd-iac.
+Model Attribute Sentiment Positional Counting Object Avg
+Qwen2VL-7B 10.0 / 28.5  2.5 / 8.5 3.5 / 20.5 6.0 / 30.5  8.0 / 50.0  6.0 / 27.6
++InstDesign 14.5 / 34.5  2.5 / 13.0 1.5 / 26.0 5.5 / 39.0 25.0 / 60.0  9.8 / 34.5
++CoT  5.0 / 15.5  6.0 / 23.5 8.5 / 30.2 6.5 / 17.0 40.5 / 59.0 13.3 / 29.0
++FewShot  3.0 / 17.0  0.5 / 9.0 1.5 / 14.5 5.0 / 29.0  2.0 / 37.0  2.4 / 21.3
++Ours 10.0 / 34.4 11.0 /16.5 14.0 / 28.3 5.0 / 37.4 51.5 / 64.0 18.4 / 36.1
+OneVision-7B 11.5/ 20.5  1.5 / 5.0  1.5 / 16.5  6.5 / 28.5 11.0 / 52.0  6.4 / 24.5
++InstDesign 16.0 / 27.0  5.5 / 12.5  6.0 / 31.5 13.5 / 30.5 34.0 / 61.5 15.0 / 32.6
++CoT 17.3 / 28.4  6.2 / 12.9  7.8 / 33.2 13.8 / 30.9 34.5 / 62.1 15.9 / 33.1
++FewShot 17.0 / 28.0  6.0 / 13.0  7.2 / 32.8 13.9 / 31.0 34.8 / 62.3 16.2 / 33.4
++Ours 19.6 / 30.5  7.8 / 13.5 10.3 / 36.4 15.1 / 29.8 35.6 / 63.5 17.7 / 34.7
+5.3 Experiments
+We verify the effectiveness of the method in controlling modality preference on MC2 and downstream tasks across Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-8B.
+We consider widely used training-free approaches as baselines: MLLM refers to employing MLLM to directly reason in modality-conflicting contexts; InstDesign uses instructions to guide modality preference direction; CoT enables complex reasoning through intermediate steps; and FewShot uses four examples to guide the models.
+For detailed implementation and results, refer to Appendix D.
+As shown in Table 2, the proposed method consistently outperforms the baseline approaches on MC2 across both settings, demonstrating its effectiveness in adjusting modality preference.
+We further assess the effectiveness of the proposed method on two types of downstream tasks: 1) multimodal machine translation (MMT) using AmbigCaps (Li et al., 2021), and 2) visual understanding on PhD (Liu et al., 2024c).
+The latter includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
+In MMT, the task should primarily ground the source-language text while treating visual information as auxiliary.
+Accordingly, we adjust modality preference toward the text modality.
+As shown in Table 2, our method yields an improvement of 1.33 BLEU score over the baselines.
+By contrast, for PhD, we steer the modality preference toward the vision modality to ground the image.
+The results in Table 3 demonstrate that our approach achieves substantial improvements across both MLLMs.
+In particular, when applied to Qwen2VL-7B, our method surpasses the best baseline by an average margin of 6.1 percentage points.
 6 Conclusion
@@
 We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference.
-We observe that all 18 tested MLLMs exhibit overall modality bias, and modality preference directions are encoded in MLLMs’ representation space.
+Besides, we find the direction of modality preference can be captured within the latent representations of MLLMs.
 Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference.
-Experiments show that the proposed method generalizes well to downstream tasks, such as hallucination mitigation and multimodal machine translation.
+Experiments show that the proposed method generalizes well to downstream tasks, such as multimodal machine translation and multimodal understanding tasks.
 Appendices
@@
 Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
-A.3 Ethical Considerations
-All images and generated text contexts, which we use to construct MC2, strictly follow guidelines designed to exclude any harmful, unethical, or offensive content.
-Human annotators are explicitly instructed to refrain from producing any personally identifiable information or inappropriate language during the annotation process.
-Furthermore, our benchmark does not involve any comparisons of harmful, ethical, or offensive content between image pairs.
+A.3 Use of LLM
+In this work, we use the LLMs including GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a), and MLLMs including LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023) to help annotate the text context in MC2, as detailed in Section B.
+We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024).
+Besides, we also utilize the LLMs to correct the grammatical errors.
 A.4 License of Assets
-All images in MC2 are publicly available from COCO [26].
+All images in MC2 are publicly available from COCO (Lin et al., 2014).
 We release our benchmark under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
@@
 To ensure reproducibility and transparency, we include the exact prompts used in our data generation process.
-These prompts were designed to generate the candidate textual contexts and corresponding answers using GPT-4o-mini [21] and DeepSeekV3 [28].
+These prompts were designed to generate the candidate textual contexts and corresponding answers using GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a).
 Below, we provide representative examples of the prompts used during dataset construction given the caption of an image, question, the answer for the question based on image and the task type for the question.
@@
 Caption: {caption} Question: {question} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
-Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer.
-The context should support the new answer, include the caption while maintaining logical consistency within the context and don’t include the final answer.
-Only one alternative answer should be generated.
-Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
 Conflict Context Generation for other tasks using GPT-4o-mini Instruction: # Given a caption of an image and a corresponding counting question with its answer, you are required to generate a single text context that provides an indirect premise leading to a new answer that fluctuates by 1 or 2 from the original answer.
@@
 Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
+Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer.
+The context should support the new answer, include the caption while maintaining logical consistency within the context and don’t include the final answer.
+Only one alternative answer should be generated.
+Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
 Human Verification
-Although the text contexts and answers generated by strong LLMs—filtered through judge MLLMs such as LLaVA1.5-7B [29] and QwenVL-7B [6]—generally yield reliable results, we further incorporate manual inspection to ensure the high quality of data annotations.
+Although the text contexts and answers generated by strong LLMs—filtered through judge MLLMs such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023)—generally yield reliable results, we further incorporate manual inspection to ensure the high quality of data annotations.
 Specifically, we verify that the visual and textual contexts are indeed in conflict, and that each modality independently supports the corresponding answer to the given question.
@@
 First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question.
-This step is particularly important because the original VQA answers in the TDIUC [22] dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
+This step is particularly important because the original VQA answers in the TDIUC (Kafle & Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
   • Conflict Verification.
@@
 In addition, we conduct random spot-checks throughout the dataset to ensure the consistency and reliability of the annotations.
+Table 4: Average text context length across different task types in the MC2 dataset.
+statistics Sport Attribute Sentiment Positional Counting Color Activity Object Avg
+Text Length 52.48 33.50 39.69 31.53 37.12 31.15 49.68 39.71 39.36
 B.2 Data Statistics
@@
 The highlighted areas indicate the points of conflict between visual and textual contexts.
-statistics Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-Text Length 52.48 33.50 39.69 31.53 37.12 31.15 49.68 39.71 39.36
-Table 5: Average text context length across different task types in the MC2 dataset.
 Appendix C Model Evaluatiion
 C.1 Evaluatiion Detail for Modality Preference
-We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B [29], LLama3.2-11B-Vision-Instruct [13], LLaVA-OneVision-7B/72B [24], CogVLM2-19B [17], mPLUG-Owl3-24-07 [49], Qwen2VL-7B [43], GLM-4V-9B [10], SPHINX-V2-1K [27], InternVL3-9B/14B/38B/78B [2], LLaVA-next-7B/13B/34B [29] and Qwen2.5VL-7B/32B/72B [7].
+We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), LLaVA-OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025).
 All the open-source models are evaluated using NVIDIA A100 or A800 GPUs.
-We also evaluate the proprietary model, GPT-4o-mini [21] via the official API.
+We also evaluate the proprietary model, GPT-4o-mini (Hurst et al., 2024a) via the official API.
 Details of single-modality context evaluation
@@
 Details of results for modality preference evaluation
-We provide the results of modality preference for several models in Table 1 in the main text.
-More detailed modality preference evaluation results are presented in Table 7.
-We evaluated 18 models, including 17 open-source models and the commercial model, GPT4o-mini.
-C.2 The Details for Controlling Modality Preference
-Noisy image and text grammar error generation
-For noisy images, we initially adopte two approaches: adding Gaussian noise to the image and applying a uniform linear motion blur kernel.
-In our experiments, both methods yield similar results in terms of model performance, indicating comparable effects on the evaluation.
-We report the results based on the Gaussian noise method in this paper.
-To generate grammatically incorrect text, we employ DeepSeek-V3 to rewrite the original text context.
-We find that DeepSeekV3 performs well in introducing localized grammatical errors while preserving the overall structure.
-We use two different prompting strategies, both of which produced similar results.
-In this paper, we report the results based on the V1 version of grammar error generation.
-The specific prompt used is shown below:
-Text Context Grammar Error Generation (V1) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes.
-These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words.
-Despite these errors, the core meaning of the sentence must remain unchanged.
-The text to be rewritten is: {original text context}.
-Output the revised text enclosed in <context2> </context2> tag.
-Only output the revised text and don’t explain anything.
-Text Context Grammar Error Generation (V2) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes.
-These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words.
-Despite these errors, the core meaning of the sentence must remain unchanged.
-Additionally, do not alter the meaning related to the provided critical word.
-The text to be rewritten is: {original text context}.
-the critical word is: {answer based on text context}.
-Output the revised text enclosed in <context1> </context1> tag.
-Only output the revised text and don’t explain anything.
-The detail for controlling modality preference using noise data
-We first test the model’s ability to understand a single noisy modality (either a noisy image context or a text context with grammar errors).
-As shown in Table 13 and Table 12, the model’s understanding of the single-modality content remains close to 95% for the noisy image context and close to 98% for the noisy text context, which is nearly the same as its performance on the original single-modality content (shown in Table 11 and Table 10).
-This rules out the possibility that the shift in modality preference due to the added noise is caused by a decrease in the model’s understanding of the single modality.
-Instead, it suggests that MLLMs may implicitly associate modality preference with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning.
-We provide the results of modality preference using either the noisy image context or the text context with grammer error in Figure 4 in main text.
-We also provide more results for more models in Table 14 and Table 15.
+We provide the results of modality preference for several models in the left panel of Figure 2 in the main text. More detailed modality preference evaluation results are presented in Table 6.
+C.2 Can the Vision Ratio provide guidance for downstream task performance?
+We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multimodal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022).
+We compute the average score on all datasets, where MME score is normalized between 0-1, as shown in Table 5.
+Model MMMU MME MMBench RealworldQA MMStar HallBench InfoVQA ChartQA Avg Vision Ratio
+Qwen2.5VL-7B 58.6 83.8 83.5 68.5 63.9 52.9 82.6 87.3 75.5 59.6
+Qwen2.5VL-72B 70.2 87.4 88.6 75.7 70.8 55.2 87.3 89.5 81.4 78.6
+InternVL3-9B 57.7 84.7 83.4 70.5 66.3 51.2 79.6 86.2 75.5 41.2
+InternVL3-14B 67.1 88.5 85.6 70.7 68.8 55.1 83.6 87.3 78.8 55.0
+InternVL3-38B 70.1 90.1 87.6 75.6 71.5 57.1 85.0 89.2 81.3 62.4
+InternVL3-78B 72.2 91.1 89.0 78.0 72.5 59.1 86.5 89.7 82.7 81.5
+LLaVA-OneVision-7B 47.9 71.2 83.2 66.3 61.7 31.6 68.8 80.0 68.4 26.3
+LLaVA-OneVision-72B 55.7 80.8 85.8 71.9 65.8 49.0 74.9 83.7 74.1 30.1
+Qwen2VL-7B 54.1 83.1 83.0 70.1 60.7 50.6 76.5 83.0 72.9 16.3
+LLaVA-Next-7B 37.6 63.2 69.2 57.8 37.6 27.6 31.6 51.9 49.8 8.5
+LLaVA-Next-13B 37.3 62.3 70.0 57.6 40.4 31.8 34.9 59.0 51.6 9.7
+LLaVA-1.5-7B 35.7 64.6 69.2 54.8 33.1 27.6 22.4 17.8 42.5 13.4
+LLaVA-1.5-13B 37.0 63.6 66.5 55.3 34.3 24.5 24.9 18.5 42.9 15.0
+Table 5: Performance comparison across benchmarks for different models measured by accuracy (%) and Vision Ratiio score (%).
+C.3 The Details for Controlling Modality Preference
 More results for controlling modality preference through instruction design.
-In Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B and Qwen2.5VL-7B.
+In the left panel of Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B.
 We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 7 and Table 8.
-For each setting, we report the probability of responses exhibiting a text preference (measured by vision-based accuracy) and the probability of responses exhibiting a vision preference (measured by text-based accuracy).
-C.3 The details of PCA Analysis
-In Section 4.2, we mentioned the PCA analysis details regarding the "Modality Preference Direction in Representation Space." Here, we provide a more in-depth description of the setup.
+For each setting, we report the results measured by $S_{vision}$, vision-based accuracy and $S_{text}$, text-based accuracy.
+The details of PCA Analysis
+In Section 4.4, we use the PCA analysis regarding the Modality Preference Direction in Representation Space.
+Here, we provide a more detailed description of the setup.
 We extract the model’s hidden representations from the last token of the input across different layers.
@@
 In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B.
-As shown in Figure 11, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-14B.
+As shown in Figure 10, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-8B.
 For each model, we selected layers with large absolute differences and small standard deviations.
 This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
-Figure 11: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B models from left to right.
-D.2 More Results for Steering Models Toward Vision or Text Modality Preference
-In the main text, we present the results of steering Qwen2VL-7B and Qwen2.5VL-7B towards vision and text modality preferences, respectively.
-Here, we provide additional results on modality preference control, including experiments with InternVL3-14B and LLaVA-OneVision-7B.
-As shown in Table 6, the proposed method significantly outperforms the baseline in both directions of modality preference.
-D.3 Downstream Task Applications
-By steering the model’s modality preference towards the vision modality, we reduce vision hallucinations in MLLMs.
-Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation.
-This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output.
-Details of Vision Hallucination Mitigation
-PhD [30] is a multimodal hallucination evaluation benchmark that increases the risk of hallucination by introducing "hitem" items, intentionally guiding the model to produce hallucinated words, thereby increasing the difficulty of the model’s responses.
-Phd-icc and Phd-ica are subsets of PhD, which respectively increase hallucination risks by introducing incorrect text context and inaccurate text context.
+Figure 10: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
+D.2 Evaluation of visual understanding and multimodal multimodal machine translation
+PhD (Liu et al., 2024c) is a visual understanding benchmark and includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
 For testing convenience, we randomly selected 1,000 samples from the original Phd-cc and Phd-ica datasets for evaluation.
-By steering modality preference towards the vision context, the model is encouraged to focus more on visual information, which helps reduce vision hallucinations.
-As shown in Table 3, the proposed method outperforms most baseline methods across nearly all subcategories of the Phd-icc and Phd-ica datasets.
-Furthermore, a detailed analysis of the two datasets reveals that most samples in Phd-iac introduce only irrelevant text context, whereas Phd-icc introduces incorrect text context, which makes it easier for the model to generate hallucinations.
-Proposed method significantly outperforms the baseline methods on Phd-icc, indicating that the proposed method is more effective in mitigating hallucinations.
-Details of Multimodal Multimodal machine translation
-Ambigcaps [25] benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets.
+By steering the model’s modality preference toward the vision modality, we strengthen its visual understanding ability and mitigate vision hallucinations in MLLMs.
+Ambigcaps (Li et al., 2021) benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets.
 We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality.
-We also use our method to steer the modality preference towards the text modality.
-As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish.
-Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-Vision MLLM-only 32.0 21.6 2.8 24.8 2.4 30.0 11.6 42.4 20.9
-Inst Design 55.6 31.2 12.0 30.8 3.2 36.4 22.4 61.2 31.6
-ours 75.6 40.8 18.0 38.0 10.0 53.2 32.0 74.4 42.8
-Text MLLM-only 8.8 48.4 60.0 52.0 47.2 21.2 52.8 9.6 37.5
-Inst Design 44.8 68.4 82.4 54.8 43.2 50.0 58.8 19.6 52.8
-ours 49.6 71.2 85.2 58.0 47.6 48.4 68.4 58.4 57.1
-Table 6: Performance by steering LLaVA-OneVision-7B towards vision and InternVL3-7B towards text.
+Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multimodal machine translation.
+This adjustment prevents the model from over-relying on visual content and from introducing spurious objects or extraneous details into the translation output.
 Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-LLaMAVision 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
+LLaMAVision-11B 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
 LLaVA1.5-7B 20.0/59.6 8.0/88.0 2.0/86.0 8.8/75.2 1.2/96.0 10.8/82.0 9.6/78.8 35.2/52.0 11.9/77.2
@@
 Owl3-24-07 60.8/31.6 16.4/72.4 10.8/85.6 22.0/69.6 8.4/88.0 28.4/60.4 17.2/71.2 60.0/29.6 28.0/63.5
-QwenVL-7B 35.2/43.6 15.2/63.2 11.2/78.0 16.4/45.6 2.8/73.6 34.0/40.8 19.2/61.6 35.2/31.2 21.1/54.7
 Qwen2VL-7B 26.4/58.0 12.4/82.8 0.8/95.6 13.2/80.4 4.0/93.6 16.0/78.8 11.6/83.6 38.0/54.0 15.3/78.3
 Qwen2.5VL-7B 65.6/12.8 45.2/46.0 18.0/68.8 46.4/38.0 51.6/39.6 70.8/20.0 42.0/43.6 77.6/14.0 52.2/35.4
-InternLM-XC2.5-7B 72.0/19.6 26.4/63.2 50.1/36.4 22.0/65.2 9.2/83.2 50.4/38.0 47.6/41.2 68.8/21.6 43.4/46.1
 GLM-4V-9B 42.0/42.4 32.4/59.2 8.8/81.6 28.0/62.4 15.2/74.4 56.8/32.8 23.3/66.0 54.0/32.8 32.6/56.5
@@
 CogVLM2-19B 53.6/29.6 28.4/47.6 10.4/56.8 17.6/56.8 6.0/62.0 36.0/35.2 34.0/39.6 67.2/19.6 31.6/43.4
-InternLM-XC2.5-7B 79.2/16.0 31.2/56.8 60.4/29.2 26.0/62.8 11.6/75.6 52.8/29.2 60.8/29.6 72.0/18.0 49.2/39.7
 GLM-4V-9B 53.2/32.8 30.4/61.2 6.0/85.6 23.2/68.0 20.0/70.0 52.4/35.6 28.0/60.8 68.0/22.0 35.2/54.5
@@
 CogVLM2-19B 98.0 95.2 99.2 96.0 94.8 98.4 98.0 99.6 97.4
-InternLM-XC2.5-7B 96.4 97.6 99.2 99.2 98.0 98.0 98.8 98.4 98.2
 GLM-4V-9B 98.4 95.2 99.6 97.2 98.8 98.4 99.6 99.6 98.4
@@
 Owl3 99.2 94.0 94.0 97.2 88.4 96.8 97.2 99.2 95.8
-QwenVL-7B 100.0 98.8 97.6 100.0 99.6 100.0 100.0 98.4 99.3
 Qwen2VL-7B 99.6 98.8 95.6 98.4 96.4 100.0 99.6 98.4 98.3
@@
 CogVLM2-19B 99.6 99.2 91.2 96.8 91.6 98.8 98.4 98.8 96.8
-InternLM-XC2.5-7B 98.8 98.8 94.0 98.0 94.8 99.6 100.0 98.8 97.9
 GLM-4V-9B 99.6 99.2 98.0 99.2 97.6 100.0 99.2 99.6 99.1
@@
 Table 10: Accuracy of question answering in the MC2 dataset when only unimodal visual context is provided.
-Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-OneVision-7B 98.0 95.2 100.0 97.2 97.2 97.2 98.0 99.2 97.8
-Qwen2.5VL-7B 98.8 96.4 100.0 99.2 97.6 97.6 98.8 99.2 98.4
-CogVLM2-19B 98.0 94.4 99.2 96.0 91.2 97.6 97.6 99.6 96.7
-InternLM-XC2.5-7B 97.6 97.2 99.2 99.2 97.6 97.6 98.8 98.8 98.2
-GLM-4V-9B 98.8 96.4 99.6 97.2 98.8 98.4 99.6 100.0 98.6
-SPHINX-V2-1K 98.4 97.6 99.2 98.8 97.6 98.4 98.4 99.6 98.5
-LLaVA1.6-7B 98.0 90.8 98.8 96.8 96.8 97.2 96.0 96.4 96.4
-Table 12: Accuracy of question answering in the MC2 dataset when only unimodal noisy textual context is provided.
-Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-OneVision-7B 98.4 95.6 94.0 98.4 78.4 99.2 95.6 98.0 94.7
-Qwen2.5VL-7B 98.8 96.4 100.0 99.2 99.2 100.0 100.0 98.8 99.3
-CogVLM2-19B 99.6 97.6 87.6 94.4 91.6 98.8 97.2 98.48 95.7
-InternLM-XC2.5-7B 98.4 96.0 94.4 94.0 95.2 100.0 97.6 95.2 96.4
-GLM-4V-9B 99.2 98.4 95.6 98.4 98.4 99.6 100.0 99.2 98.6
-SPHINX-V2-1K 98.4 97.2 90.0 98.8 98.4 97.6 96.8 98.0 96.9
-LLaVA1.6-7B 98.4 94.8 92.4 93.6 84.0 96.8 93.6 95.6 93.7
-Table 13: Accuracy of question answering in the MC2 dataset when only unimodal noisy visual context is provided.
-Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-OneVision-7B 42/29.2 31.2/41.6 4.4/88 31.6/44 4.4/82 39.2/29.6 16.4/66.8 50.8/26 27.5/50.9
-Qwen2.5VL-7B 74.0/12.0 53.20/36.40 24.4/62.4 58/29.2 54.8/35.2 78/16 52.8/35.6 82.4/12.4 59.7/29.9
-InternLM-XC2.5-7B 75.6/18.4 32.8/57.2 56.0/36.0 26.4/59.2 11.6/78.0 61.2/28.0 52.8/36.4 74.0/20.0 48.8/41.7
-GLM-4V-9B 48.0/31.6 37.6/48.8 10.0/81.2 34.0/56.4 18.8/70.8 64.8/27.2 33.6/55.6 61.6/25.2 38.6/49.6
-SPHINX-V2-1K 46.0/44.8 18.0/77.6 1.2/98.0 21.6/70.8 11.6/80.8 30.0/64.8 30.8/60.0 62.8/28.8 27.8/65.7
-CogVLM2-19B 46.0/39.6 34.0/47.2 12.4/68.4 22.8/52.0 8.4/70.0 37.2/39.2 30.0/55.2 62.8/24.4 31.7/49.5
-LLaVA1.6-7B 11.2/71.2 6.4/83.2 1.2/91.6 5.2/76.4 0.4/88.8 7.2/71.6 5.6/70.4 28.8/44.4 8.3/74.7
-Table 14: Accuracy of question answering in the MC2 dataset when both noisy textual context and visual context are provided but the instruction does not specify which modality context should be used.
-Values are reported as vision-based accuracy/text-based accuracy for each model.
-Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
-OneVision-7B 24/44 16.4/66.4 2/96.8 17.2/66.4 0.80/87.2 20.4/47.2 8/78.8 35.6/42.8 15.5/66.2
-Qwen2.5VL-7B 48.8/23.6 25.6/64.8 6/88 33.2/59.6 39.2/54.8 58.8/32.8 30/56.8 68/56.8 38.7/50.3
-InternLM-XC2.5-7B 52.8/33.6 15.2/76.8 24.4/64.4 12.8/78.4 3.6/90.4 33.2/56.8 24.4/64.8 54.8/37.2 27.7/62.8
-GLM-4V-9B 32.4/50.4 19.6/73.2 4.0/91.6 21.2/69.2 11.2/78.4 41.6/46.8 15.2/76.8 41.2/42.0 23.3/66.1
-SPHINX-V2-1K 29.2/58.0 11.2/86.4 0.8/98.8 16.8/80.4 5.6/90.4 18.0/76.4 18.0/76.0 52.8/40.4 19.1/75.9
-CogVLM2-19B 32.0/45.2 22.0/60.0 3.6/84.8 14.0/60.8 6.4/73.2 18.8/49.2 18.8/65.6 46.0/36.4 20.2/59.4
-LLaVA1.6-7B 8.4/79.2 4.0/87.2 0.4/94.8 3.2/82.4 0.4/92.0 4.4/77.6 4.4/76.8 22.8/50.8 6.0/80.1
-Table 15: Accuracy of question answering in the MC2 dataset when both textual and noisy visual contexts are provided but the instruction does not specify which modality context should be used.
-Values are reported as vision-based accuracy/text-based accuracy for each model.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. 2.
-    Limitations
-  5. Question: Does the paper discuss the limitations of the work performed by the authors?
-  6. Answer: [Yes] See discussions in Suppl.
-  7. 3.
-    Theory assumptions and proofs
-  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  9. Answer: [N/A]
-  10. 4.
-    Experimental result reproducibility
-  11.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  12. Answer: [Yes] See details in Suppl.
-  13. 5.
-    Open access to data and code
-  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  15. Answer: [Yes] See details in Suppl.
-  16. 6.
-    Experimental setting/details
-  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  18. Answer: [Yes] See details in Suppl.
-  19. 7.
-    Experiment statistical significance
-  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  21. Answer: [Yes] See details in Suppl.
-  22. 8.
-    Experiments compute resources
-  23. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  24. Answer: [Yes] See details in Suppl.
-  25. 9.
-    Code of ethics
-  26. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  27. Answer: [Yes]
-  28. 10.
-    Broader impacts
-  29. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  30. Answer: [Yes]
-  31. 11.
-    Safeguards
-  32.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  33. Answer: [Yes] See details in Suppl.
-  34. 12.
-    Licenses for existing assets
-  35.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  36. Answer: [Yes]
-  37. 13.
-    New assets
-  38. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  39. Answer: [Yes] See detail in Suppl.
-  40. 14.
-    Crowdsourcing and research with human subjects
-  41.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  42. Answer: [Yes] See details in Suppl.
-  43. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  44.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  45. Answer: [Yes] See details in Suppl.
-  46. 16.
-    Declaration of LLM usage
-  47.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  48. Answer: [Yes] See details in Suppl.

### v3

@@
-Evaluating and Steering Modality Preferences in Multimodal Large Language Model
+Evaluating and Steering Modality Preferences in Multi-modal LLMs
 Abstract
-Multimodal large language models (MLLMs) have achieved remarkable success on complex multimodal tasks.
-However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multimodal contexts.
+Multi-modal large language models (MLLMs) have achieved remarkable success on complex multi-modal tasks.
+However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multi-modal contexts.
 To study this question, we introduce MC2 benchmark, which constructs controlled evidence-conflict scenarios to systematically evaluate modality preference in decision-making.
-Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performances of MLLMs.
+Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performance of MLLMs.
 Further analysis shows that modality preference can be controlled by instruction guidance and captured within the latent representations of MLLMs.
 Built on these insights, we propose a probing and steering method based on representation engineering to explicitly control modality preference without requiring additional fine-tuning.
-This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple downstream applications, including multimodal visual understanding and multimodal machine translation.
+This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple multi-modal understanding and reasoning tasks.
 footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
+Machine Learning, ICML
 1 Introduction
-Multimodal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video).
-Recent advances demonstrate their exceptional capabilities on complex tasks with multimodal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multimodal dialogue systems (Sun et al., 2022).
-Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multimodal inputs.
+multi-modal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video).
+Recent advances demonstrate their exceptional capabilities on complex tasks with multi-modal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multi-modal dialogue systems (Sun et al., 2022).
+Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multi-modal inputs.
+Figure 1: Illustrations of evaluating modality preference. Using multi-modal conflict context pairs to evaluate modality preference.
 To investigate this, one line of work (Fu et al., 2024; Amara et al., 2024) compares model performance on unimodal input, providing either only text or only image input for the same question.
-Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes of the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu & Frank, 2024; 2022).
-However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multimodal scenarios, where information from different modalities naturally co-occur.
+Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes in the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu and Frank, 2024, 2022).
+However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multi-modal scenarios, where information from different modalities naturally co-occurs.
 In this paper, we provide a controllable setup to study the modality preference in MLLMs.
-As shown in the left panel of Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting evidence from different modalities.
-In this way, we can determine the modality preference based on the answer given by MLLMs.
-To dismiss the effect of confounding factors including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we annotate and select perception-level tasks that demonstrate accurate question comprehension and reliable single-modality recognition.
-Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight perception-level tasks with 2,000 carefully selected samples.
+As shown in the Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting pieces of evidence from different modalities.
+In this way, we can determine modality preference based on the answer given by MLLMs.
+To enable a rigorous and fair assessment, we eliminate existing confounding factors including the internal knowledge of MLLMs and inherent understanding capabilities.
+Consequently, we curate and annotate the dataset to ensure that the model achieves precise question and single-modality context comprehension.
+Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight tasks with 2,000 carefully selected samples.
 Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 20 representative MLLMs.
 Our study reveals several intriguing findings:
-  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the right panel of Figure 1, and modality preference can serve as a useful indicator of downstream task performances.
+  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the Figure 2, and modality preference can serve as a useful indicator of downstream task performance.
   • Internal attention patterns toward specific modalities give rise to modality preference, and the underlying factors can be traced to the training data recipe and the scale of the MLLMs.
@@
 Built on these, we propose a modality preference probing and steering method based on representation engineering (Zou et al., 2023) to explicitly amplify the modality preference without additional fine-tuning.
-Experimental results show that the proposed method leads to notable performance improvements on multimodal visual understanding and multimodal machine translation.
+Experimental results show that the proposed method leads to notable performance improvements on multi-modal understanding and reasoning tasks.
 Our main contributions are summarized as follows:
-Figure 1: Illustrations of evaluating modality preference.
-Left: Using multimodal conflict context pairs to evaluate modality preference.
-Right: Quantified scores for Vision and Text modalities, where a higher score indicates a stronger preference toward the corresponding modality.
-Other represents irrelevant predictions, which are discarded during the statistics.
   • We introduce MC2 to comprehensively evaluate modality preferences in MLLMs and highlight the significance of modality preference in correlating the downstream task performance.
-  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representation, providing insights into multimodal reasoning.
+  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representations, providing insights into multi-modal reasoning.
   • We propose a training-free method that steers modality preference via representation-level intervention, enabling controllable preference adjustment and enhancing performance on downstream tasks.
 2 Related Work
+2.1 multi-modal Large Language Models
+In recent years, Large Language Models (LLMs) have achieved impressive results across a wide spectrum of tasks (Lu et al., 2022; Zhang et al., 2024b; Guo et al., 2024; Xu et al., 2025; Huang and Chang, 2023).
+Building upon these, multi-modal Large Language Models (MLLMs) have extended such capabilities, demonstrating strong performance in visual question answering, structured information understanding, and visual navigation (Zhu et al., 2024; Zhang et al., 2025a; Liu and Liu, 2025; Wei et al., 2023; Zhang et al., 2025b).
+Recent research has turned to systematically evaluating their biases and failure modes (J.
+Liu, Y.
+Fu, R.
+Xie, R.
+Xie, X.
+Sun, F.
+Lian, Z.
+Kang, and X.
+Li (2024c); 11; K.
+Amara, L.
+Klein, C.
+Lüth, P.
+Jäger, H.
+Strobelt, and M.
+El-Assady (2024); A.
+Alishahi, G.
+Chrupała, and T.
+Linzen (2019); M.
+Chen, Y.
+Cao, Y.
+Zhang, and C.
+Lu (2024b)), aiming to better understand their limitations and guide the development of more reliable models.
 2.2 Modality Preference
@@
 Data factors influencing modality preference.
-Research on data-related factors (Guo et al., 2023; Chen et al., 2024; Leng et al., 2024) explores how properties of multimodal datasets give rise to and reinforce modality preference.
-In particular, many samples in multimodal datasets can be resolved correctly by relying on information from only a single modality.
-When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024).
+Research on data-related factors (Guo et al., 2023; Chen et al., 2024a; Leng et al., 2024) explores how properties of multi-modal datasets give rise to and reinforce modality preference.
+In particular, many samples in multi-modal datasets can be resolved correctly by relying on information from only a single modality.
+When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024a).
 Furthermore, their inclusion in evaluation benchmarks artificially inflates performance metrics, as models can exploit these unimodal shortcuts instead of performing genuine cross-modal integration (Leng et al., 2024; Winterbottom et al., 2020).
@@
 Evaluating the intrinsic modality preference in MLLMs.
-Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multimodal models optimize for multimodal inputs.
-Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024).
-While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multimodal systems.
-Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu & Frank, 2022; 2024).
-However, in real-world multimodal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining truly modality preference.
-Wu et al.
-(2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts.
-However, conflict detection is only one facet of multimodal reasoning and does not comprehensively reflect a model’s modality preference when processing multimodal contexts.
-In this work, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts.
-Compared to prior work, we carefully control confounding variables such as input quality, question-understanding ability, and internal model knowledge, and construct a modality context conflict dataset, enabling a more rigorous evaluation of modality preference and uncovering new insights.
-Furthermore, we design a flexible method which can controllably steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
+Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multi-modal models optimize for multi-modal inputs.
+Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024a).
+While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multi-modal systems.
+Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu and Frank, 2022, 2024).
+However, in real-world multi-modal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining true modality preference.
+(Wu et al., 2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multi-modal contexts.
+However, conflict detection is only one facet of multi-modal reasoning and does not comprehensively reflect a model’s modality preference when processing multi-modal contexts.
+In this work, we investigate the behavioral tendencies of MLLMs by examining their responses to questions within scenarios involving conflicting multi-modal contexts.
+Unlike prior studies, we decouple modality preference from confounding variables—such as internal model knowledge and varying understanding capabilities—by ensuring the models can accurately comprehend the provided context.
+Furthermore, we design a flexible method to steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
 2.3 Representation Engineering
-Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023).
+Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023b; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023).
 Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs (Greenblatt et al., 2023; Stolfo et al., 2024; Wu et al., 2024; Xu et al., 2024; Zou et al., 2023).
-In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs).
-Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
+In this work, we extend this paradigm to a novel setting: controlling modality preference in multi-modal large language models (MLLMs).
+Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, allowing for flexible and precise modulation of model behavior.
 3 The MC2 Benchmark
@@
 3.1 Data Design Philosophy
-To reveal the modality preference of MLLMs when processing truly multimodal inputs, we construct modality context conflict pairs and carefully control various confounding factors that may affect the measurement of modality preference, such as the internal knowledge of MLLMs, question comprehension, and the understanding of unimodal contexts.
-Therefore, we focus on evaluating perception-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs.
-Besides, we ensure that the evaluated models can accurately answer questions based solely on the context from a single modality, minimizing the influence of confounding factors.
+Modality preference is a fundamental behavioral tendency to favor one modality over another, irrespective of MLLMs’s internal knowledge and inherent understanding capabilities.
+To enable a rigorous and fair assessment, we eliminate these confounding factors by ensuring that models can accurately comprehend the multi-modal context and respond based solely on the provided information.
+We elaborate on this design choice below: 1) Neutralizing Knowledge Bias: By ensuring models can derive answers directly from the context, we decouple modality preference from the interference of varying internal knowledge bases.
+2) Establishing a Level Playing Field: By verifying that all models can comprehend the multi-modal context, we establish a “common ground”—a baseline of competence that ensures a fair comparison across MLLMs with diverse capabilities.
+Finally, we construct modality-conflict pairs to evaluate and compare the modality preferences of different MLLMs in a controlled environment.
+Figure 2: Results of modality preference across different MLLMs.
+Left: Quantified scores for Vision, Others and Text modalities using $S_{\text{vision}}$, $S_{\text{others}}$ and $S_{\text{text}}$ as well as Vision Ratio.
+Right: Trends of Vision Ratio with respect to model parameter size.
 3.2 Semi-automated Data Construction Pipeline
-In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design philosophy outlined in Section 3.1.
-The dataset is derived from the TDIUC (Kafle & Kanan, 2017) dataset, sourced from MS-COCO (Lin et al., 2014), widely adopted in model development to ensure the evaluated models can recognize the images.
-We select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the vision context and the image caption capcap for each sample from TDIUC as the foundation for data annotation.
-The pipeline follows these steps:
+In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous process to ensure the reliability of the dataset.
+The data is derived from the TDIUC (Kafle and Kanan, 2017), sourced from MS-COCO (Lin et al., 2014).
+Therefore, the data is widely adopted in model development, ensuring that most MLLMs can recognize the images.
+Firstly, we select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the image and the caption capcap for each sample from TDIUC.
+Then the pipeline follows these steps:
 Textual Context Construction.
-Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with the $c^{v}$ and capcap in terms of overall scene semantics.
+Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with $c^{v}$ and capcap in terms of overall scene semantics.
 We prompt DeepSeekV3 (Liu et al., 2024a) and ChatGPT4o-mini (Hurst et al., 2024b) to generate a distractor answer $A^{t}$ to $q$, together with $c^{t}$ that plausibly supports $A^{t}$, using carefully crafted instructions.
 For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection.
-To ensure that all evaluated MLLMs demonstrate strong recognition capabilities for both visual and textual contexts, we employ several basic MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
+To ensure that all evaluated MLLMs can correctly recognize both visual and textual contexts, we employ several weak MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
 Human Verification.
@@
 Iterative Refinement.
-The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which helps identify and rectify potential errors, thereby enhancing the dataset quality.
-Modality Context Conflict Dataset. To this end, we construct MC2, a modality context conflict dataset including 2000 samples.
-The detailed instruction templates for textual context generation, the detailed manual annotation procedures, the data annotation format along with sample cases and dataset statistics are provided in Appendix B.
+The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which rectifies potential errors, thus enhancing the data quality.
+MC2.
+To this end, we develop MC2, a modality-conflict dataset comprising 2,000 samples spanning diverse tasks.
+The detailed construction procedures for textual context generation, the manual annotation, the data format and dataset statistics are provided in Appendix B.
 3.3 Question Design and Evaluation Metric
@@
 We reformulate the original questions with ChatGPT-4o-mini (Hurst et al., 2024a) into a binary-choice format.
-To further reduce potential position bias in multimodal inputs, we adopt a consistent evaluation strategy, similar to Liu et al.
+To further reduce potential position bias in multi-modal inputs, we adopt a consistent evaluation strategy, similar to Liu et al.
 (2024d).
@@
 Evaluation Metric.
-Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this idea to evaluate the modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities.
-More importantly, through the careful design of our benchmark, we establish as a basis that the model can reliably understand both modalities in isolation.
-As shown in Table 9 and Table 10 in Appendix, all models achieve over 95% accuracy when provided with either textual or visual context.
+Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this to evaluate modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities.
+More importantly, through the careful design of the benchmark, we establish as a basis that the model can reliably understand both modalities in isolation.
+As shown in Table 16 and 17 in Appendix, all MLLMs achieve over 95% accuracy, provided with either textual or visual context.
 Building on this, our metric evaluates modality preference by assessing how the model’s responses align with textual or visual input when the two provide conflicting signals.
-The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the responses are ambiguous, uncertain, or inconsistent with either modality, which are discarded from further analysis.
+The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the response is uncertain, or inconsistent with either modality, which is discarded from further analysis.
 Then, we naturally define the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/({S_{\text{vision}}+S_{\text{text}}})$, where $S_{\text{vision}}$ or $S_{\text{text}}$ denotes the score of the vision or text modality, computed as the proportion of samples whose responses are categorized as Vision or Text across the dataset.
-$S_{others}$ is the proportion of samples whose responses are categorized as Others.
+$S_{\text{others}}$ is the proportion of samples whose responses are categorized as Others.
 Vision Ratio greater than 0.5 indicates that the model tends to favor visual context over text.
+Figure 3: Analysis of modality preferences.
+Left: Trends of Vision Ratio and Attention Ratio.
+Middle: Vision Ratio with Respect to the proportion of multi-modal conflict-context training data and different MLLMs.
+Right: Relationship between visual understanding ability (quantified as the average accuracy across seven widely used benchmarks) and the modality preference (measured by Vision Ratio).
 4 Modality Preferences in MLLMs
@@
 4) Can modality preference be controlled?
-This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
-Figure 2: Results of modality preference across different MLLMs.
-Left: Quantified scores for Vision, Others Text modalities using $S_{vision}$, $S_{others}$ and $S_{text}$ as well as Vision Ratio.
-Right: Trends of Vision Ratio with respect to model parameter size across different MLLMs.
+This investigation helps uncover the underlying mechanisms of modality preference and leverage these insights to downstream tasks.
 4.1 Which Modality Do MLLMs Prioritize?
-We use the MC2 benchmark to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
+We use MC2 to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
 Different MLLMs exhibit different modality preferences.
-As described in Section 3.3, we quantify modality preference using Vision Ratio, with the results presented in the left panel of Figure 2 and detailed in Table 6.
+As described in Section 3.3, we quantify modality preference using Vision Ratio in the left panel of Figure 2 and detailed in Table 13.
 We observe that all MLLMs exhibit clear modality preference, with most models showing a strong preference for text; for instance, LLaVA1.5-7B attains only a 13.4% Vision Ratio.
-This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu & Frank, 2024; Wu et al., 2025).
-Interestingly, the Qwen2.5VL and InternVL3 show a certain degree of preference towards the vision modality.
-Larger MLLMs exhibit stronger preferences for the vision modality.
+This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu and Frank, 2024; Wu et al., 2025).
+Interestingly, Qwen2.5VL and InternVL3 show a certain degree of visual preference.
+Larger MLLMs exhibit stronger visual preferences.
 We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference.
 As shown in the right panel of Figure 2, we observe that for all model families, the preference for the vision modality increases with the model size.
-And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases.
-However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their sizes increase.
+And Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases.
+However, other models maintain a noticeable textual preference as model sizes increase.
+To validate the reliability of the evaluation, we conduct a sensitivity analysis for the sample number of MC2 in Appendix E.1.
 Vision Ratio aligns with human preference.
 We further verify whether the Vision Ratio can serve as a human-level measure of modality preference in MLLMs.
-We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, LLaVA-OneVision-7B, InternVL3-14B, and LLaVA1.5-7B.
+We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, OneVision-7B, InternVL3-14B, and LLaVA1.5-7B.
 In addition, we craft prompts to elicit explicit reasoning chains from the models, specifically targeting their reliance on visual or textual information.
 To ensure labeling reliability, three expert annotators independently annotate the expressed modality preference for each response, with the final label determined by majority vote.
-The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with the ones given by human ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%.
+The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with those given by human annotators ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%.
 This indicates that the Vision Ratio can act as a reliable, automated proxy for human assessment of modality preference.
-Figure 3: Analysis of modality preferences.
-Left: Trends of Vision Ratio and multimodal Attention Ratio across different models.
-Middle: Vision Ratio with Respect to the Proportion of Multimodal Conflict-Context Training Data and different MLLMs.
-Right: Relationship between visual understanding ability, quantified as the average accuracy across seven widely used benchmarks and the modality preference measured by Vision Ratio.
 4.2 What Factors Drive These Preferences?
-Given that different MLLMs display varying modality preferences, we examine two primary sources of such differences: the internal attention distribution and the training factors.
+Given that different MLLMs display varying modality preferences, we examine two primary sources: the internal attention distribution and the training factors.
 Different allocation of attention across modalities.
 We compute the mean attention scores over all token positions from both modalities and define the ratio of visual attention to total attention as the Attention Ratio.
-By analyzing Qwen2.5VL-7B and LLaVA-OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3.
+By analyzing Qwen2.5VL-7B and OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3.
 This alignment suggests that MLLMs distribute attention unevenly between modalities, which in turn contributes to their divergent modality preferences.
-Impact of model scale and training data recipe.
-Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM.
-Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multimodal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference.
-2) Larger LLMs are more capable of shifting their preference during training;
-To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multimodal conflict contexts, adjusting their preferences toward text or vision.
-We then measured changes with the Vision Ratio.
-We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio.
-As shown in the middle panel of Figure 3, increasing the proportion of multimodal contexts consistently leads to larger preference shifts, supporting Hypothesis 1.
-This suggests that multimodal inputs maybe create more challenging training conditions, leading to stronger shifts of preference.
-Furthermore, Qwen2.5VL-7B exhibites greater shifts than Qwen2.5VL-3B under the same conditions, supporting Hypothesis 2.
-This indicates that larger LLMs demonstrate stronger learning ability and adapt more effectively.
-4.3 Can the Vision Ratio provide guidance for downstream task performance?
-To demonstrate the correlation between the modality preference and performance for downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs.
-Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in the Appendix C.2.
-We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3.
-The results reveal a strong positive association between the Vision Ratio and visual understanding ability across the evaluated MLLMs.
-Specifically, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that the Vision Ratio provides a highly reliable indicator of visual understanding task performance.
-4.4 Can Modality Preference Be Controlled?
-We employ instruction guidance to investigate whether modality preference can be controlled, and conduct a latent space representation analysis to examine the mechanisms, underlying the preference adjustment.
-Details of the experimental design and results are provided in Appendix C.3.
 Figure 4: Analysis of modality preference under instruction-guidance.
@@
 Middle: Representation shifts under instruction-guided interventions.
-Right: layer-wise absolute difference and standard deviation of hidden states between different instruction.
+Right: Layer-wise absolute difference and standard deviation of hidden states between different instructions.
+Impact of model scale and training data recipe.
+Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM.
+Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multi-modal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference; 2) Larger LLMs are more capable of shifting their preference during training.
+To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multi-modal conflict contexts to adjust preferences toward text or vision.
+We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio.
+As shown in the middle panel of Figure 3, increasing the proportion of multi-modal contexts consistently leads to larger preference shifts, supporting Hypothesis 1.
+This suggests that multi-modal inputs may create more challenging conditions, leading to greater preference shifts.
+Furthermore, Qwen2.5VL-7B exhibits greater shifts than Qwen2.5VL-3B under the same training data, supporting Hypothesis 2.
+This indicates that larger LLMs demonstrate stronger adaptability.
+4.3 Can the Vision Ratio provide guidance for downstream task performance?
+As a foundational behavioral prior, the identified modality preference can inform how a model integrates information across modalities.
+As such, the findings can offer relevant insights into the model’s behavior in deeper cross-modal understanding tasks.
+To demonstrate the correlation between the modality preference and performance on downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs.
+Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in Appendix C.2.
+We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3.
+The results reveal a strong positive association between Vision Ratio and visual understanding ability across all MLLMs.
+Notably, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that Vision Ratio provides a highly reliable indicator of visual understanding ability.
+4.4 Can Modality Preference Be Controlled?
+We investigate whether modality preference can be controlled, and conduct latent space representation analysis to examine the underlying mechanisms.
+The detailed setting and results are provided in Appendix C.3.
 Modality preference can be guided through instruction design.
 We investigate the impact of instruction design on modality preference, for instance, by explicitly directing the model to rely on a specific modality when answering a question.
-Specifically, we evaluate the modality preference measured by Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text or vision preference instruction (Inst-text, Inst-vision).
-As illustrated in the left panel of Figure 4, instructions that steer the model toward a particular modality effectively shape its modality preference.
+Specifically, we evaluate the modality preference using Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text and vision preference instruction (Inst-text and Inst-vision).
+As shown in the left panel of Figure 4, instructions, probing the models toward any modality effectively shape their modality preferences.
 Modality preference direction in representation space.
 To further understand how the intervention methods influence modality preference internally, we analyze the hidden representations of the models.
-Specifically, we apply Principal Component Analysis (PCA; Abdi & Williams, 2010) to the hidden states to identify the dominant direction corresponding to modality preference shifts.
-The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, aligning with the modality specified by the instruction.
-The PCA direction further reveals that the model’s internal states are sensitive to modality control cues, which motivates us to develop representation techniques for adjusting modality preference and to apply these insights to downstream tasks in Section 5.
+Specifically, we apply Principal Component Analysis (PCA; Abdi and Williams, 2010) to the hidden states to identify the dominant direction of modality preference.
+The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, revealing that the model’s internal states are sensitive to modality control.
+This motivates us to develop specific techniques to adjust modality preference in Section 5.
 Figure 5: Overall framework of the proposed method.
@@
 Modality Preference Steering selects the target layer during the second inference and adds the scaled modality preference direction to the representation at the corresponding layer at each inference step.
-5 Representation Based Modality Preference Steering
-Inspired by the representation behavior discussed in Section 4.4, we propose to use the representation engineering (Zou et al., 2023) to steer the modality preference, controlling the model’s behavioral expression.
-As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
+5 Method
+Inspired by the behavior discussed in Section 4.4, we propose to use representation engineering (Zou et al., 2023) to steer modality preference.
+As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§6).
 5.1 Modality Preference Probing
@@
 Inspired by the pre-training Next Token Prediction objective of decoder-only MLLMs (Hurst et al., 2024a) and the method to extract classification features (Feucht et al., 2024), we collect neural activity from the last token in the input text.
-The process involves probing modality preference through two pairs of requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’).
-Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$.
-Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states on the two queries at the last token of the input at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM.
-We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs.
-More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $l$ for a given query as:
+The process involves probing modality preference through two requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’).
+Let us denote these inputs by $q^{v}$ and $q^{t}$, and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$.
+Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states for the two queries at the last token at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM.
+We identify the direction of modality preference by computing the difference in the hidden states between the paired inputs.
+Formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $\ell$ for a given query as:
 $\mathbf{u}^{t}_{\ell}=\frac{1}{N}\sum_{i}^{N}\big(\mathbf{x}_{i,\ell}^{t}-\mathbf{x}_{i,\ell}^{v}\big).$ (1)
+Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality, and Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{\text{vision}}$ and $S_{\text{text}}$.
+Preference Model MLLM InstDesign CoT FewShot Ours
+[…]
+InternVL3-8B 20.9 31.6 36.7 38.2 42.8
+[…]
+OneVision-7B 37.5 52.8 53.1 49.8 57.1
 Averaging over different queries allows us to capture the activation values most closely associated with modality preference, independent of questions.
 As shown in the right panel of Figure 4, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples.
-We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers.
-Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 5.2.
+We observe that layers 20–23 exhibit higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers.
+Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 6.
 Similar patterns are observed for Qwen2VL-7B, Qwen2.5VL-7B, LLaVA-OneVision and InternVL3, as detailed in Appendix D.1.
-5.2 Modality Preference Steering
-After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$.
-The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior.
-In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference.
-Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states.
-This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities.
-Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
-$\mathbf{s}^{t}_{\ell}=w\mathbf{u}^{t}_{\ell},\ \text{where}\ w=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\mathbf{x}^{t}_{i,\ell}\|}{\|\mathbf{u}^{t}_{\ell}\|}$ (2)
-Finally, during the second inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward the text modality.
-Similarly, steering models towards vision modality is performed towards the opposite direction.
-In the final implementation, no additional data or labels are introduced.
-We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the modality preference.
-Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality and steering Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{vision}$ and $S_{text}$.
-Preference Model MLLM InstDesign CoT FewShot Ours
-[…]
-OneVision-7B 20.9 31.6 36.7 38.2 42.8
-[…]
-InternVL3-8B 37.5 52.8 53.1 49.8 57.1
-Table 2: Multimodal translation results for Ambigcaps (Li et al., 2021). BLEU scores are reported for English (En) $\leftrightarrow$ Turkish (Tr).
-Method En->Tr Tr->En
-Qwen2.5VL-7B 8.92 18.56
-      +Inst towards vision  8.21 (-0.71) 16.09 (-2.47)
-      +Inst towards text  9.45 (+0.53) 18.98 (+0.42)
-      +Ours 10.22 (+1.30) 19.89 (+1.33)
+Table 2: Performance of the proposed method on the reasoning tasks (MathVista, TallyQA and VSR) measured by accuracy for Qwen2.5VL-7B and multi-modal translation task (Ambigcaps) measured by BLEU scores for English (En) $\leftrightarrow$ Turkish (Tr).
+Dataset CoT InstDesign Ours
+MathVista 50.0 59.0 62.3
+TallyQA 61.6 74.4 76.6
+VSR 45.6 51.3 53.2
+Ambigcaps (En→Tr) 9.03 9.45 10.2
+Ambigcaps (Tr→En) 18.63 18.98 19.89
 Table 3: Performance of the proposed method on the visual understanding benchmark, Phd (Liu et al., 2024c). we report the accuracy results on the phd-icc/phd-iac.
@@
 +Ours 19.6 / 30.5  7.8 / 13.5 10.3 / 36.4 15.1 / 29.8 35.6 / 63.5 17.7 / 34.7
+5.2 Modality Preference Steering
+Figure 6: The attention scores of the generated token toward the vision and text contexts at the 22th layer using InstDesign (Left) and the proposed method (Right).
+After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w$.
+The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior.
+In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference.
+Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states.
+This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities.
+Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
+$\mathbf{s}^{t}_{\ell}=w\mathbf{u}^{t}_{\ell},\ \text{where}\ w=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\mathbf{x}^{t}_{i,\ell}\|}{\|\mathbf{u}^{t}_{\ell}\|}$ (2)
+Finally, during the second inference, we adjust the hidden states at the selected layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward text modality.
+Similarly, steering towards vision modality is performed towards the opposite direction and no additional data or labels are introduced.
 5.3 Experiments
-We verify the effectiveness of the method in controlling modality preference on MC2 and downstream tasks across Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-8B.
-We consider widely used training-free approaches as baselines: MLLM refers to employing MLLM to directly reason in modality-conflicting contexts; InstDesign uses instructions to guide modality preference direction; CoT enables complex reasoning through intermediate steps; and FewShot uses four examples to guide the models.
-For detailed implementation and results, refer to Appendix D.
-As shown in Table 2, the proposed method consistently outperforms the baseline approaches on MC2 across both settings, demonstrating its effectiveness in adjusting modality preference.
-We further assess the effectiveness of the proposed method on two types of downstream tasks: 1) multimodal machine translation (MMT) using AmbigCaps (Li et al., 2021), and 2) visual understanding on PhD (Liu et al., 2024c).
-The latter includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
-In MMT, the task should primarily ground the source-language text while treating visual information as auxiliary.
-Accordingly, we adjust modality preference toward the text modality.
-As shown in Table 2, our method yields an improvement of 1.33 BLEU score over the baselines.
-By contrast, for PhD, we steer the modality preference toward the vision modality to ground the image.
-The results in Table 3 demonstrate that our approach achieves substantial improvements across both MLLMs.
-In particular, when applied to Qwen2VL-7B, our method surpasses the best baseline by an average margin of 6.1 percentage points.
+We evaluate the proposed method for controlling modality preference on MC2 and multiple downstream tasks across four representative MLLMs: Qwen2VL-7B, Qwen2.5VL-7B, OneVision-7B, and InternVL3-8B.
+We compare our approach against several widely-used training-free baselines: MLLM (zero-shot direct response), InstDesign (instruction-based guidance for modality preference), CoT (utilizing intermediate processing steps), and FewShot (four-shot in-context learning).
+Implementation details are provided in Appendix D.
+As shown in Table 1, our method consistently outperforms all baselines on MC2 across both evaluation settings.
+These results demonstrate the superior effectiveness of our approach in precisely adjusting modality preference compared to standard prompting techniques.
+We further assess our method across a variety of downstream scenarios: 1) Multi-modal reasoning tasks on MathVista (Lu et al., 2023), TallyQA (Acharya et al., 2019) and VSR (Liu et al., 2023a); 2) Multi-modal understanding tasks on PhD (Liu et al., 2024c) and AmbigCaps (Li et al., 2021), a multi-modal machine translation (MMT) task.
+The Phd dataset includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
+As shown in Table 2, our method consistently surpasses both the CoT and InstDesign baselines.
+We observe that by steering the model toward a visual preference, our method increases reliance on the original visual input.
+This effectively prevents the final decision from being misled by potential vision hallucination that often emerges during the sequential steps of CoT.
+On the AmbigCaps translation task, our method achieves a significant improvement of $1.33$ BLEU score over the strongest baseline.
+Furthermore, Table 3 demonstrates substantial improvements in multi-modal understanding.
+Specifically, when applied to Qwen2VL-7B, our method surpasses the best-performing baseline by an average margin of $6.1$ percentage points.
+Additional results for more tasks and models are available in Appendix E.2.
+5.4 In-depth Analysis for Steering Method
+To investigate the internal mechanism of the steering method, we analyze the attention scores of the generated token toward the vision and text contexts using the proposed steering method and the InstDesign method with the samples from $MC^{2}$.
+In Figure 6, we visualize the attention distribution at the 24th layer by steering the Qwen2.5VL-7B at the 22nd layer towards the text modality using the case in Figure 12, as detailed in Appendix E.4.
+We observe that after applying the steering method, the model’s attention weight toward the text modality significantly increases.
+This change clearly demonstrates that the steering mechanism successfully alters the modality preference by enhancing the model’s dependency on the text modality.
+More ablation experiments, the analysis of the latency, memory usage and the prerequisites for the proposed method are detailed in Appendix E.3 and  E.4.
 6 Conclusion
-This paper investigates modality preference in multimodal large language models (MLLMs).
+This paper investigates modality preference in multi-modal large language models (MLLMs).
 We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference.
@@
 Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference.
-Experiments show that the proposed method generalizes well to downstream tasks, such as multimodal machine translation and multimodal understanding tasks.
-Appendices
-All codes, data, and instructions for our MC2 can be found in https://anonymous.4open.science/r/Modality-Preference-8016. MC2 is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
-Our supplementary materials are summarized as follows:
+Experiments show that the proposed method generalizes well to downstream tasks, such as multi-modal reasoning and understanding tasks.
+Impact Statement
+The proposed MC2 evaluates the modality preference of MLLMs.
+Understanding which modality a model prioritizes could be used to circumvent safety mechanisms (e.g., hiding harmful content in the favored modality), making it harder for filters to detect inappropriate content.
+Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
 Appendix A Discussion
 A.1 Limitations
-This paper investigates the modality preference in multimodal large language models (MLLMs) using a controlled experiment setup with a modality conflict dataset.
+This paper investigates the modality preference in multi-modal large language models (MLLMs) using a controlled experiment setup with a modality conflict dataset.
 In constructing the dataset, we employs LLaVA1.5-7B and QwenVL-7B to filter samples and ensure that most models could answer questions correctly based on a single modality.
@@
 Therefore, devising a more efficient and elegant method for sample selection may be of greater importance.
-A.2 Social Impacts
-The proposed MC2 evaluates the modality preference of MLLMs.
-Understanding which modality a model prioritizes could be used to circumvent safety mechanisms (e.g., hiding harmful content in the favored modality), making it harder for filters to detect inappropriate content.
-Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
+A.2 Strategic Insights for Future Research
+Based on the experiment analysis, we highlight two key principles for designing effective multi-modal learning strategies: 1) As observed in Section 4.2, increasing the proportion of data with multi-modal context (TVQA) accelerates the adjustment of modality preference.
+Therefore, for MLLMs—especially those inheriting parameters from LLMs—introducing a more TVQA data can more rapidly mitigate the inherent language bias, facilitating faster convergence toward balanced multi-modal learning.
+2) We identify that the Vision Ratio serves as a reliable indicator of visual understanding, and adjusting modality preference improves downstream performance.
+Consequently, combined with the first point, selecting an appropriate ratio of TVQA training data can adjust modality preferences to satisfy the specific requirements of different downstream.
 A.3 Use of LLM
 In this work, we use the LLMs including GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a), and MLLMs including LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023) to help annotate the text context in MC2, as detailed in Section B.
-We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024).
+We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024).
 Besides, we also utilize the LLMs to correct the grammatical errors.
@@
 Only one alternative answer should be generated.
-Caption: {caption} Question: {question} Answer: {answer based on vision context} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
+Caption: {caption} Question: {question} Answer: {answer based on vision context} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
 Conflict Context Generation of for other tasks using DeepSeekV3 Instruction: # Given the caption of an image and a corresponding {task-type} type question with its answer, now you are required to generate a text context as a premise that supports a new distractor answer for the question.
@@
 Only one alternative answer should be generated.
-Caption: {caption} Question: {question} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
+Caption: {caption} Question: {question} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
 Conflict Context Generation for other tasks using GPT-4o-mini Instruction: # Given a caption of an image and a corresponding counting question with its answer, you are required to generate a single text context that provides an indirect premise leading to a new answer that fluctuates by 1 or 2 from the original answer.
@@
 Only one alternative answer should be generated.
-Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
+Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
 Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer.
@@
 Only one alternative answer should be generated.
-Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
+Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
 Human Verification
@@
 First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question.
-This step is particularly important because the original VQA answers in the TDIUC (Kafle & Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
+This step is particularly important because the original VQA answers in the TDIUC (Kafle and Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
   • Conflict Verification.
@@
 C.1 Evaluatiion Detail for Modality Preference
-We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), LLaVA-OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025).
+We assess open-source multi-modal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025).
 All the open-source models are evaluated using NVIDIA A100 or A800 GPUs.
@@
 C.2 Can the Vision Ratio provide guidance for downstream task performance?
-We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multimodal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022).
+We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multi-modal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022).
 We compute the average score on all datasets, where MME score is normalized between 0-1, as shown in Table 5.
@@
 InternVL3-78B 72.2 91.1 89.0 78.0 72.5 59.1 86.5 89.7 82.7 81.5
-LLaVA-OneVision-7B 47.9 71.2 83.2 66.3 61.7 31.6 68.8 80.0 68.4 26.3
+OneVision-7B 47.9 71.2 83.2 66.3 61.7 31.6 68.8 80.0 68.4 26.3
 LLaVA-OneVision-72B 55.7 80.8 85.8 71.9 65.8 49.0 74.9 83.7 74.1 30.1
@@
 More results for controlling modality preference through instruction design.
-In the left panel of Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B.
+In the left panel of Figure 4, we provide the Vision Ratio results for OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B.
 We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 14 and Table 15.
-For each setting, we report the results measured by $S_{vision}$, vision-based accuracy and $S_{text}$, text-based accuracy.
+For each setting, we report the results measured by $S_{\text{vision}}$, vision-based accuracy and $S_{\text{text}}$, text-based accuracy.
 The details of PCA Analysis
@@
 In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B.
-As shown in Figure 10, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-8B.
+As shown in Figure 11, we present the visualization of hidden states for OneVision-7B, Qwen2VL-7B, and InternVL3-8B.
 For each model, we selected layers with large absolute differences and small standard deviations.
 This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
-Figure 10: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
-D.2 Evaluation of visual understanding and multimodal multimodal machine translation
+Figure 11: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
+D.2 Evaluation of visual understanding and multi-modal multi-modal machine translation
 PhD (Liu et al., 2024c) is a visual understanding benchmark and includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
@@
 Ambigcaps (Li et al., 2021) benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets.
-We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality.
-Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multimodal machine translation.
+We evaluate the multi-modal machine translation (MMT) task on this dataset using the Qwen2.5VL-7B model.
+multi-modal contexts in MMT are both complementary and contradictory: the visual information provides helpful context for translation, but the potential for conflicting, non-visual signals can interfere with grounding the source language.
+Consequently, the proposed method is designed to steer the modality preference toward the text modality to ensure robustness against these visual-textual conflicts.
+Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multi-modal machine translation.
 This adjustment prevents the model from over-relying on visual content and from introducing spurious objects or extraneous details into the translation output.
+Appendix E More Experiment Analysis
+E.1 The Sensitivity Analysis for the Evaluation of Modality Preference
+To verify whether the current 2k-sample scale of $MC^{2}$ is sufficient to ensure the stability of both the preference evaluation and steering results, we conduct a sensitivity analysis by randomly selecting a specified quantity of samples from each category for assessment.
+As shown in Table 6, we calculate the Vision Ratio for OneVision-7B and Qwen2.5VL-7B.
+Results demonstrate that as the sample size increases per category, the Vision Ratio begins to stabilize around 150 samples.
+These experiments suggest that the current dataset size for each task is sufficient to ensure the stability of evaluation of modality preference.
+In the future, we would like to expand MC2 with a wider variety of tasks (e.g., texture recognition) and modalities (e.g., Audio), further enhancing its comprehensiveness and generalizability.
+Table 6: The sensitivity analysis for the evaluation of modality preference. We  randomly select a specified quantity of samples from each category for assessment, measured by Vision Ratio.
+Model 25 50 75 100 125 150 175 200 225 250
+OneVision-7B 29.2 28.0 29.1 29.6 28.1 26.4 26.6 26.7 26.7 26.3
+Qwen2.5VL-7B 56.7 59.4 60.2 59.9 60.4 59.8 60.2 59.7 59.8 59.6
+E.2 More Results for Downstream Tasks
+E.2.1 More results for Phd dataset
+We evaluate the performance of Phd dataset using the proposed method for LLaVA-1.6-7B and the results are in Table 7.
+We observe that LLaVA-1.6-7B, due to its severe text preference, only achieves an average ACC score of $0.7$ on the PHD-icc subset.
+Applying our method significantly boosts the model’s performance across the two subsets, thereby demonstrating the cross-model generalization of our approach.
+Table 7: Performance of the proposed method on the visual understanding benchmark, Phd using LLaVA-1.6-7B.
+Phd Method Attribute Sentiment Positional Counting Object Avg
+Phd-icc LLaVA-1.6-7B 0.5 0.0 0.0 1.5 1.5 0.7
+InstDesign 2.0 0.5 0.5 1.5 9.5 2.8
+Ours 3.5 1.5 1.0 3.0 15.0 4.8
+Phd-iac LLaVA-1.6-7B 5.5 8.0 4.0 10.5 29.0 11.4
+InstDesign 7.5 12.5 14.5 15.5 44.5 18.9
+Ours 11.5 20.0 20.5 22.0 51.0 25.0
+E.3 Ablation Study
+In this section, we conduct the detailed ablation study to analyze the proposed method.
+E.3.1 The number of probing samples
+We compute the preference direction in Equation 1 using varied sample sizes but test the steering performance on the complete $MC^{2}$ dataset.
+We report the $S_{\text{Vision}}$ and $S_{\text{Text}}$ results for OneVision-7B and Qwen2.5VL-7B in Table 8.
+We observe that steering performance remains stable even when the steering vector is derived from a limited number of samples.
+Table 8: The ablation study for the varied sample number of computing the preference direction.
+Model 25 50 75 100 125 150 175 200 225 250
+OneVision-7B 56.0 55.9 56.3 57.1 57.0 57.3 57.3 57.4 57.2 57.1
+Qwen2.5VL-7B 62.2 62.2 61.7 62.6 64.0 62.8 62.7 63.1 62.8 63.6
+E.3.2 The diversity of probing samples
+To study the impact of data diversity for probing task, we experiment by using only a single task for probing and applying the resulting vector to steer all other tasks.
+We report the $S_{\text{Vision}}$ and $S_{\text{Text}}$ for OneVision-7B and Qwen2.5VL-7B for entire dataset in Table 9.
+We observe that OneVision-7B achieves competitive performance compared to our initial implementation for using nearly each probing task, with Qwen2.5VL-7B showing similar success on over half the tasks.
+Further analysis finds that the most effective single-probing tasks are those where the initial modality preference change was more pronounced.
+Table 9: The ablation study for the varied diversities of computing the preference direction.
+Model Sport Attribute Sentiment Positional Counting Color Activity Object Ours
+OneVision-7B 58.6 55.8 59.3 57.8 59.0 53.4 58.5 58.6 57.1
+Qwen2.5VL-7B 64.1 59.5 70.0 51.9 47.4 65.5 63.0 60.6 63.6
+E.3.3 Different steering intensities
+To investigate the performance with varied steering intensities, we introduce a scaling coefficient $\lambda$ to the steering weight $w$ in Equation 2 to change the steering intensity and conduct a test on $MC^{2}$ for both OneVision-7B and Qwen2.5-VL-7B, reporting the $S_{\text{Vision}}$ or $S_{\text{Text}}$ scores in Table 10.
+We observe that performance drops for both models with decreased steering intensity, indicating insufficient steering.
+As intensity increases, the performance of OneVision-7B significantly drops, exhibiting clear over-steering at $\lambda=2.0$ which leads to destruction of language capabilities.
+Conversely, for Qwen2.5VL-7B, the steering effect continues to enhance up to $\lambda=1.75$, only significantly degrading beyond $\lambda=2.0$, demonstrating a wider safe steering margin in its representation space.
+Table 10: The ablation study for the varied steering intensities.
+Model 0.125 0.25 0.5 0.75 1.0 1.25 1.5 1.75 2.0 2.25 2.5 2.75 3.0
+OneVision-7B 40.2 45.5 51.8 52.6 57.1 56.5 32.6 12.1 3.9 0.1 0.0 0.0 0.0
+Qwen2.5VL-7B 40.8 44.1 49.4 53.2 63.6 72.3 76.0 70.7 45.8 15.2 16.2 12.5 7.8
+<image> represents the visual input displayed below
+User: $\langle\text{image}\rangle$ Text Context: The table was adorned with a vibrant bouquet of flowers and a charming ceramic sheep, while the surrounding chairs, crafted from smooth, polished wood, complemented the rustic yet elegant setting.
+In case of inconsistency, prioritize the text context.
+Question: What are the chairs made of? A. wicker    B. wood
+Assistant: $\langle\text{output}\rangle$
+Vision-based answer: wicker
+Text-based answer: wood
+Baseline answer: A. wicker ✗
+Proposed steering answer: B. wood ✓
+Figure 12: Example of a cross-modal conflict sample from our diagnostic dataset.
+The highlighted text indicates the ground-truth modality specified by the instruction, which the proposed steering method successfully follows while the baseline fails.
+Figure 13: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 23rd layer.
+Figure 14: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 24th layer.
+Figure 15: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 25th layer.
+Figure 16: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 26th layer.
+Figure 17: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 27th layer.
+Figure 18: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 28th layer.
+E.4 Details for the application of the proposed method
+E.4.1 Details of in-depth analysis for steering method
+We provide the detailed description of the case for in-depth analysis for steering method in Figure 6 in Section 5.4.
+Besides, we also provide more attention analysis for the case in the different layers in Figure 1314151617 and 18.
+We observe that across all subsequent layers following steering modality preference towards text at the 21 th layer, our method significantly increases the model’s attention weight toward the text modality, surpassing the corresponding vision attention weight.
+This change clearly demonstrates that the steering mechanism successfully alters the modality preference by enhancing the model’s dependency on the text modality.
+E.4.2 Latency and Memory
+The proposed method consists of two phases, probing and steering.
+The probing phase is conducted offline and the resulting steering vector is cached for reuse.
+During the actual steering phase, we simply load this cached steering vector, which incurs minimal memory overhead and does not add meaningful computational cost to the inference process.
+We measure the single-sample inference latency (seconds) (without Flash-Attention acceleration and batch inference) for our method compared to the MLLM baseline (MLLM-only) in Table 11.
+The results show that the steering phase introduces negligible latency compared to the MLLM-only baseline, and the overhead is confined to the initial offline probing stage.
+Furthermore, all three methods require nearly identical memory requirements.
+E.4.3 The prerequisites for implementation of the proposed method
+Based on Representation Engineering (Greenblatt et al., 2023; Xu et al., 2024), the proposed method requires capturing an explicit modality preference direction vector to realize behavioral adjustment.
+The approach succeeds when such a vector can be reliably extracted, as seen in models like Qwen2.5VL-7B.
+However, the method fails in cases such as LLaVA-1.5-7B, primarily due to its limited ability to follow instructions for preference adjustment, which prevents the capture of a meaningful direction vector.
+Table 11: The comparison of inference time between MLLM-only and the proposed method including probing and steering phases.
+Dataset MC2 Phd-icc
+MLLM-only 1.99 1.84
+Probing (Offline) 2.21 2.12
+Steering 2.00 1.84
+Besides, we observe a localized performance drop in the Attribute subset of the Phd-icc benchmark in Table 3.
+We attribute this to the limitation of applying a single global steering vector, which may fail to accommodate instance-level granularity—where fine-grained, sample-specific features are required for optimal alignment.
+Despite these isolated cases, the overall benchmark performance improves, underscoring the effectiveness of our approach, especially considering that it requires no external data or fine-tuning.
+E.5 The Further Experiment without modality preference prior
+Our current approach intentionally select the steering direction based on known task requirements.
+This design has been proven to be pragmatic and effective for many real-world applications where the optimal modality is clear.
+In addition, our method can be readily integrated with a training-free priority detection method to enable dynamic preference selection.
+To demonstrate this, we conduct the following experiment:
+Dataset Construction: We modify $MC^{2}$ dataset by degrading the quality of one modality context so that only one modality is reliable, and the ground-truth answer aligns with it.
+We use QA accuracy to measure model performance on this new dataset.
+Task Design: 1) Each sample requires a specific reliable modality. 2) All samples share the same reliable modality in a task. Each task contains 200 samples.
+Solution: We apply a causal analysis approach (Parcalabescu and Frank, 2024) to identify the reliable modality.
+For each sample, we first measure the change of predicted answer probability when removing either the image or the text context.
+The larger the drop, the more important that modality is for the given sample.
+For Task1, we determine the reliable modality for a specific sample by comparing the probability drops.
+For Task2, by aggregating the reliable modalities across all samples via majority voting, we determine the preferred modality for the specific task.
+Results: For the identification of reliable modality, we achieve an accuracy of $85.3\%$ for all samples in Task1; we reach $100\%$ accuracy for task-level identification in Task2 (thus, performance on Task 2 is equivalent to knowing the steering preference in advance).
+Next, we evaluate the performance of the proposed method on Task1, measured by QA accuracy in Table 12.
+The results show that steering with predicted preference yields significant gains over base models and closely matches the performance of the “preference prior” setting.
+This confirms that our method can be simply adapted to autonomously prioritize modalities based on input quality or task needs.
+Table 12: The performance of the proposed method on the revised $MC^{2}$ without modality preference prior measured by Accuracy.
+Method Task1
+OneVision-only 25.4
++Steering with preference prior 40.7
++Steering with predicted preference 37.5
+Qwen2.5VL-7B-only 49.1
++Steering with preference prior 62.7
++Steering with predicted preference 58.2
+Table 13: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided but the instruction does not specify which modality context should be used.
+Values are reported as vision-based accuracy/text-based accuracy for each model.
 Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
@@
 GPT-4o-mini 94.4/3.2 35.6/47.6 60.4/28.4 22.0/58.9 19.4/59.2 34.8/36.4 71.2/20.4 78.4/12.8 52.0/33.4
-Table 6: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided but the instruction does not specify which modality context should be used.
+Table 14: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on visual modality context.
 Values are reported as vision-based accuracy/text-based accuracy for each model.
@@
 LLaVA1.6-34B 84.8/12.0 48.0/36.4 62.8/24.0 34.0/52.0 38.8/44.4 76.4/14.4 62.0/18.4 80.4/12.4 60.9/26.8
-Table 7: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on visual modality context.
+Table 15: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on the textual modality.
 Values are reported as vision-based accuracy/text-based accuracy for each model.
@@
 InternVL3-14B 28.4/44.8 14.0/68.4 3.6/82.4 21.2/54.8 28.0/43.2 24.8/50.0 17.2/58.8 55.2/19.6 24.0/52.8
-Table 8: Accuracy of question answering in the MC2 dataset when both textual and visual contexts are provided and the instruction explicitly directs the model to answer based on the textual modality.
-Values are reported as vision-based accuracy/text-based accuracy for each model.
+Table 16: Accuracy of question answering in the MC2 dataset when only unimodal textual context is provided.
 Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
@@
 GPT-4o-mini 97.6 97.2 99.6 98.6 97.4 98.4 98.4 100.0 98.4
-Table 9: Accuracy of question answering in the MC2 dataset when only unimodal textual context is provided.
+Table 17: Accuracy of question answering in the MC2 dataset when only unimodal visual context is provided.
 Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
@@
 GPT-4o-mini 100.0 92.0 95.6 100.0 100.0 96.0 96.4 96.0 97.0
-Table 10: Accuracy of question answering in the MC2 dataset when only unimodal visual context is provided.